### v1

# MetaGFN: Exploring Distant Modes with Adapted Metadynamics for Continuous GFlowNets

Abstract
Generative Flow Networks (GFlowNets) are a class of generative models that sample objects in proportion to a specified reward function through a learned policy. They can be trained either on-policy or off-policy, needing a balance between exploration and exploitation for fast convergence to a target distribution. While exploration strategies for discrete GFlowNets have been studied, exploration in the continuous case remains to be investigated, despite the potential for novel exploration algorithms due to the local connectedness of continuous domains. Here, we introduce Adapted Metadynamics, a variant of metadynamics that can be applied to arbitrary black-box reward functions on continuous domains. We use Adapted Metadynamics as an exploration strategy for continuous GFlowNets. We show three continuous domains where the resulting algorithm, MetaGFN, accelerates convergence to the target distribution and discovers more distant reward modes than previous off-policy exploration strategies used for GFlowNets.
1 Introduction
Generative Flow Networks (GFlowNets) are a type of generative model that samples from a discrete space $\mathcal{X}$ by sequentially constructing objects via actions taken from a learned policy $P_{F}$ Bengio et al. (2021a). The policy $P_{F}(s,s^{\prime})$ specifies the probability of transitioning from some state $s$ to some other state $s^{\prime}$. The policy is parameterised and trained so that, at convergence, the probability of sampling an object $x\in\mathcal{X}$ is proportional to a specified reward function $R(x)$. GFlowNets offer advantages over more traditional sampling methods, such as Markov chain Monte Carlo (MCMC), by learning an amortised sampler, capable of single-shot generation of samples from the desired distribution. Since GFlowNets learn a parametric policy, they can generalise across states, resulting in higher performance across various tasks Bengio et al. (2021a); Malkin et al. (2022); Zhang et al. (2022); Jain et al. (2022); Deleu et al. (2022); Jain et al. (2023); Hu et al. (2023); Zhang et al. (2023); Shen et al. (2023b) and applications to conditioned molecule generation Shen et al. (2023b), maximum likelihood estimation in discrete latent variable models Hu et al. (2023), structure learning of Bayesian networks Deleu et al. (2022), scheduling computational operations Zhang et al. (2023), and discovering reticular materials for carbon capture Cipcigan et al. (2023).
Although originally conceived for discrete state spaces, GFlowNets have been extended to more general state spaces, such as entirely continuous spaces, or spaces that are hybrid discrete-continuous Lahlou et al. (2023). In the continuous setting, given the current state, the policy specifies a continuous probability distribution over subsequent states, and the probability density over states $x\in\mathcal{X}$ sampled with the policy is proportional to a reward density function $r(x)$. The continuous domain unlocks more applications for GFlowNets, such as molecular conformation sampling Volokhova et al. (2023) and continuous control problems Luo et al. (2024).
GFlowNets are trained like reinforcement learning agents. Trajectories of states are generated either on-policy or off-policy, with the terminating state $x\in\mathcal{X}$ providing a reward signal for informing a gradient step on the policy parameters. GFlowNets therefore suffer from the same training pitfalls as reinforcement learning. One such issue is slow temporal credit assignment, which has thus far been addressed by designing more effective loss functions, such as detailed balance Bengio et al. (2021b), trajectory balance Malkin et al. (2022) and sub-trajectory balance Madan et al. (2022).
Besides loss functions, another aspect of GFlowNet training is the exploration strategy for acquiring training samples. Exclusively on-policy learning is generally inadequate as it leads to inefficient exploration of new modes. More successful strategies therefore rely on off-policy exploration. For the discrete setting, numerous exploration strategies have been proposed including $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al. (2022), Thompson sampling Rector-Brooks et al. (2023) and Local Search GFlowNets Kim et al. (2024). While these approaches can be generalised to the continuous domain, there is no literature benchmarking their effectiveness in this setting.
Sampling in the continuous setting is a common occurrence in various domains such as molecular modelling Hawkins (2017); Yang et al. (2019) and Bayesian inference Shahriari et al. (2016). The local connectedness of a continuous domain allows for novel exploration strategies that are not directly applicable in the discrete setting. In this work, we create MetaGFN, an exploration algorithm for continuous GFlowNets inspired by metadynamics, an enhanced sampling method widely used for molecular modelling Laio and Parrinello (2002). The main contributions of this work are:
  • Presenting MetaGFN, an algorithm created by adapting metadynamics to black box rewards and continuous GFlowNets;
  • Proving that the proposed Adapted Metadynamics is consistent and reduces to standard metadynamics in a limit;
  • Showing empirically that MetaGFN outperforms existing GFlowNets exploration strategies.
The rest of the paper is as follows. In Section 2 we review the theory of discrete and continuous GFlowNets. We present Adapted Metadynamics and MetaGFN in Section 3. In Section 4, we evaluate MetaGFN against other exploration strategies, showing that MetaGFN outperforms existing exploration strategies in two continuous environments. We finish with limitations and conclusions in Sections 5 and 6. Code for MetaGFN is available at [link in camera-ready].
2 Preliminaries
2.1 Discrete GFlowNets
In a GFlowNet, the network refers to a directed acyclic graph (DAG), denoted as $G=(\mathcal{S},\mathcal{A})$. Nodes represent states $s\in\mathcal{S}$, and edges represent actions $s\rightarrow s^{\prime}\in\mathcal{A}$ denoting one-way transitions between states. The DAG has two distinguishable states: a unique source state $s_{0}$, that has no incoming edges, and a unique sink state $\perp$, that has no outgoing edges. The set of states, $\mathcal{X}\subset\mathcal{S}$, that are directly connected to the sink state are known as terminating states. GFlowNets learn forward transition probabilities, known as a forward policy $P_{F}(s^{\prime}|s)$, along the edges of the DAG so that the resulting marginal distribution over the terminal states complete trajectories (the terminal distribution) , denoted as $P^{\perp}(x)$, is proportional to a given reward function $R\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. GFlowNets also introduce additional learnable objects, such as a backward policy $P_{B}(s|s^{\prime})$, which is a distribution over the parents of any state of the DAG, to create losses that train the forward policy. Objective functions for GFlowNets include flow matching (FM), detailed balance (DB), trajectory balance (TB) and subtrajectory balance (STB) Bengio et al. (2021a, b); Malkin et al. (2022); Madan et al. (2022). During training, the parameters of the flow objects are updated with stochastic gradients of the objective function applied to batches of trajectories. These trajectory batches can be obtained either directly from the current forward policy or from an alternative algorithm that encourages exploration. These approaches are known as on-policy and off-policy training respectively.
2.2 Continuous GFlowNets
Continuous GFlowNets extend the generative problem to continuous spaces Lahlou et al. (2023), where the analogous quantity to the DAG is a measurable pointed graph (MPG) Nummelin (1984). MPGs can model continuous spaces (e.g., Euclidean space, spheres, tori), as well as hybrid spaces, with a mix of discrete and continuous components, as often encountered in robotics, finance, and biology Neunert et al. (2020); Swiler et al. (2012); Bortolussi and Policriti (2008).
Definition 2.1 (Measurable pointed graph (MPG)).
Let $(\bar{\mathcal{S}},\mathcal{T})$ be a topological space, where $\bar{\mathcal{S}}$ is the state space, $\mathcal{T}$ is the set of open subsets of $\bar{\mathcal{S}}$, and $\Sigma$ is the Borel $\sigma$-algebra associated with the topology of $\bar{\mathcal{S}}$. Within this space, we identify the source state $s_{0}\in\bar{\mathcal{S}}$ and sink state $\perp\in\bar{\mathcal{S}}$, both distinct and isolated from the rest of the space. On this space we define a reference transition kernel $\kappa\vcentcolon\bar{\mathcal{S}}\times\Sigma\rightarrow[0,+\infty)$ and a backward reference transition kernel $\kappa^{b}\vcentcolon\bar{\mathcal{S}}\times\Sigma\rightarrow[0,+\infty)$. The support of $\kappa(s,\cdot)$ are all open sets accessible from $s$. The support of $\kappa^{b}(s,\cdot)$ are all open sets where $s$ is accessible from. Additionally, these objects must be well-behaved in the following sense:
  • (i)
    Continuity: For all $B\in\Sigma$, the mapping $s\mapsto\kappa(s,B)$ is continuous.
  • (ii)
    No way back from the source: The backward reference kernel has zero support at the source state, i.e. for all $B\in\Sigma$, $\kappa^{b}(s_{0},B)=0$.
  • (iii)
    No way forward from the sink: When at the sink, applying the forward kernel keeps you there, i.e. $\kappa(\perp,\cdot)=\delta_{\perp}(\cdot)$, where $\delta_{\perp}$ is the Dirac measure of the sink state.
  • (iv)
    A fully-explorable space: The number of steps required to be able to reach any measurable $B\in\Sigma$ from the source state with the forward reference kernel is bounded.
The set of objects $(\bar{\mathcal{S}},\mathcal{T},\Sigma,s_{0},\perp,\kappa,\kappa^{b})$ then defines an MPG.
Note that the support of $\kappa(s,\cdot)$ and $\kappa^{b}(s,\cdot)$ are analogous to the child and parent sets of a state $s$ in a DAG. Similarly, a discrete GFlowNet’s DAG satisfies discrete versions of (ii), (iii), and (iv).
The set of terminating states $\mathcal{X}$ are the states that can transition to the sink, given by $\mathcal{X}=\left\{s\in\mathcal{S}\vcentcolon\kappa(s,\{\perp\})>0\right\}$, where $\mathcal{S}\vcentcolon=\bar{\mathcal{S}}\setminus\{s_{0}\}$. Trajectories $\tau$ are sequences of states that run from source to sink, $\tau=(s_{0},\dots,s_{n},\perp)$. The forward Markov kernel $P_{F}\vcentcolon\bar{\mathcal{S}}\times\Sigma\rightarrow[0,\infty)$ and backward Markov kernel $P_{B}\vcentcolon\bar{\mathcal{S}}\times\Sigma\rightarrow[0,\infty)$ have the same support as $\kappa(s,\cdot)$ and $\kappa^{b}(s,\cdot)$ respectively, where being a Markov kernel means states are mapped to probability measure, hence $\int_{\bar{\mathcal{S}}}P_{F}(s,ds^{\prime})=\int_{\bar{\mathcal{S}}}P_{B}(s,ds^{\prime})=1$. A flow F is a tuple $F=(f,P_{F})$, where $f:\Sigma\rightarrow[0,\infty)$ is a flow measure, satisfying $f(\{\perp\})=f(s_{0})=Z$, where $Z$ is the total flow.
The reward measure is a positive and finite measure $R$ over the terminating states $\mathcal{X}$, we denote the density of this reward measure as $r(x)$, for $x\in\mathcal{X}$. A flow $F$ is said to satisfy the reward-matching conditions if
$R(dx)=f(dx)P_{F}(x,\{\perp\}).$
If a flow satisfies the reward-matching conditions and trajectories are recursively sampled from the Markov kernel $P_{F}$ starting at $s_{0}$, the resulting measure over terminating states, $P^{\perp}(B)$, is proportional to the reward: $P^{\perp}(B)=\frac{R(B)}{R(\mathcal{X})}$ for any $B$ in the $\sigma$-algebra of terminating states Lahlou et al. (2023).
Objective functions for discrete GFlowNets generalise to continuous GFlowNets. However, in the continuous case, the forward policy $\hat{p}_{F}:\mathcal{S}\times\bar{\mathcal{S}}\rightarrow[0,\infty)$, backward policy $\hat{p}_{B}:\mathcal{S}\times\bar{\mathcal{S}}\rightarrow[0,\infty)$ and parameterised flow $\hat{f}:\mathcal{S}\rightarrow[0,\infty)$ parameterise the $P_{F}$, $P_{B}$ transition kernels and flow measure $f$ on an MPG. Discrete GFlowNets parameterise log transition probabilities and flows on a DAG. In this work, we consider DB, TB and STB losses. For a complete trajectory $\tau$, the TB loss can be written as
$L_{TB}(\tau)=\left(\log\frac{Z_{\theta}\prod_{t=0}^{n}\hat{p}_{F}(s_{t},s_{t+1};\theta)}{r(s_{n})\prod_{t=0}^{n-1}\hat{p}_{B}(s_{t+1},s_{t};\theta)}\right)^{2},$
where $Z_{\theta}$ is the parameterised total flow (see Appendix A for the DB and STB loss functions).
2.3 Exploration strategies for GFlowNets
GFlowNets can reliably learn using off-policy trajectories, a key advantage over hierarchical variational models Malkin et al. (2023). For optimal training, it is common to use a replay buffer and alternate between on-policy and off-policy (exploration) batches Shen et al. (2023a). Exploration strategies for discrete GFlowNets include $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al. (2022), Local Search GFlowNets Kim et al. (2024), and Thompson sampling (TS), which outperforms the others in grid and bit sequence domains Rector-Brooks et al. (2023). TS aims to bias exploration in regions where there is high uncertainty. When the forward policy is parameterised as an MLP, this is achieved using an ensemble of $K\in\mathbb{Z}^{+}$ policy heads with a common network torso. During training, an ensemble member is randomly sampled and used to generate a trajectory $\tau$. In a training batch, each ensemble member is included with probability $p$ and parameters are updated by taking a gradient step on the total loss of $\tau$ over all included members.
No comparative literature exists on exploration strategies for continuous GFlowNets, but many methods can be adapted, and we do so here. For example, for TS in the continuous setting, policy heads parameterise forward policy functions instead of log probabilities. Another strategy unique to the continuous setting is what we call noisy exploration. This strategy involves introducing an additive noise parameter, denoted as $\bar{\sigma}$, to the variance parameters in the forward policy distribution, where the value of $\bar{\sigma}$ is scheduled to gradually decrease to zero over the course of training.
2.4 Metadynamics
Molecular dynamics (MD) uses Langevin dynamics (LD) Pavliotis (2014), a stochastic differential equation modelling particle motion with friction and random fluctuations, to simulate atomic trajectories that ergodically sample a molecule’s Gibbs measure, $\rho_{\beta}(x)\propto e^{-\beta V(x)}$. Here, $x$ and $p$ are atomic positions and momenta, $V(x)$ is the molecular potential, and $\beta$ is thermodynamic beta. [^1] We review Langevin dynamics in Appendix B. If the potential $V(x)$ has multiple local minima, then unbiased LD can get trapped in these minima, which can lead to inefficient sampling. Metadynamics is an algorithm that enhances sampling by regularly depositing repulsive Gaussian bias potentials at the centre of an evolving LD trajectory Laio and Parrinello (2002). At time $t>0$, metadynamics modifies LD such that the conservative component of the LD force is given by the negative gradient of the total potential, i.e. $-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$, where $V_{\text{bias},t}$ is the cumulative bias at time $t$. Bias potentials are typically chosen to only vary in the direction of user-specified, low-dimensional collective variables (CVs), $z(x):\mathcal{X}\rightarrow\mathcal{Z}$, mapping from the original space $\mathcal{X}$ to the CV space $\mathcal{Z}$. For biomolecules, typical CVs include protein backbone angles or distances between charge centres; these are quantities that drive and parameterise rare-event transitions. As the bias potential progressively fills up the potential landscape, energetic barriers are reduced, thus accelerating exploration, eventually ensuring uniform diffusion in CV space. In practice, the bias is defined on a regular grid, which limits CV space dimensionality to 5-10 due to exponential memory costs. Thus, identifying good CVs is crucial for effective metadynamics simulations in applications such as drug discovery, chemistry and materials science De Vivo et al. (2016); Laio and Gervasio (2008). In this work, we adapt the original metadynamics algorithm discussed above. The method has also seen numerous extensions. For a detailed review, see Bussi and Laio (2020) and the references therein Bussi and Laio (2020).
3 MetaGFN: Adapted Metadynamics for GFlowNets
Training GFlowNets in high-dimensional continuous spaces requires exploration, especially if valuable reward peaks are separated by large regions of low reward. In some tasks, prior knowledge of the principal manifold directions along which the reward measure varies can reduce the effective dimension required for exploration. This is where exploration algorithms that guarantee uniform sampling in that low-dimensional manifold, such as metadynamics, become most effective. With this intuition in mind, we adapt metadynamics to the black-box reward setting of continuous GFlowNets.
[…]
Kernel density potential Metadynamics force computations require the gradient of the total potential, where $-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$. Using the above assumptions, we have $\nabla V=-\nabla r(x)/(\beta^{\prime}r(x))$. However, $r(x)$ is often a computationally expensive black-box function, and its gradient, $\nabla r(x)$, is unknown. While finite differences can estimate $\nabla r(x)$ for smooth, low-dimensional reward distributions, this approach is impractical in high-dimensional spaces. We avoid finite difference calculations by computing a kernel density estimate (KDE) of $V$. In particular, we assume collective variables $z(x)=(z_{1}(x),\dots,z_{k}(x))$, and compute both the KDE potential and bias potentials in the CV space. Here, each $z_{i}$ is a one-dimensional coordinate, and $z\vcentcolon\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{Z}$ is $k$-dimensional. As the potentials are stored on a low-dimensional grid, gradient computations are guaranteed to be cheap relative to the cost of evaluating $r(x)$. We use $\hat{V}(z)$ to denote the KDE estimate in CV space.
To compute $\hat{V}(z)$, we maintain two separate KDEs: $\hat{N}(z)$ for the histogram of metadynamics states and $\hat{R}(z)$ for cumulative rewards. We update these KDE estimates on the fly at the same time the bias potential is updated. If $\mathcal{Z}\cong\mathbb{R}^{k}$, we use Gaussian kernels with update rules: [^2] If $\mathcal{Z}\cong\mathbb{T}^{k}$ ($k$-torus), we use von Mises distributions instead of Gaussians.
$\hat{N}(z)_{t}=\hat{N}(z)_{t-1}+\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
$\hat{R}(z)_{t}=\hat{R}(z)_{t-1}+r(x_{t})\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
where $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\mathbb{R}^{k}$ is the kernel width, $x_{t}$ is the latest metadynamics sample and $z_{t,i}$ is the corresponding $i^{th}$ CV coordinate of $x_{t}$. The KDE potential $\hat{V}(x)$ is then computed as:
$\hat{V}(x)=-\frac{1}{\beta^{\prime}}\log\left(\frac{\hat{R}(x)}{\hat{N}(x)+\epsilon}+\epsilon\right),$
where $\epsilon>0$. We found empirically that $\epsilon$ ensured numerical stability by preventing division by zero and bounding the potential above by $\log\left(1/\epsilon\right)/\beta^{\prime}$. Taking the ratio of a reward and frequency KDE means that $\hat{V}$ rapidly and smoothly adjusts when new modes are discovered, in particular, we prove that $\hat{V}$ eventually discovers all reward modes in the CV space. More precisely,
Theorem 3.1.
If the collective variable $z(x)$ is analytic with a bounded domain, then
$\lim_{\epsilon\rightarrow 0}\left(\lim_{\sigma\rightarrow 0}\left(\lim_{t\rightarrow\infty}\hat{V}(z,t)\right)\right)=V,$ (1)
where $V=V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$ is the marginal potential in the CV space if $z(x)$ is not invertible. If $z(x)$ is invertible, $V$ is the original potential $V(x)$ in the original coordinates.
The proof is in Appendix C.
Implementation details We set the potential energy beta ($\beta^{\prime}$) equal to the Langevin dynamics beta ($\beta$). This allows a physical interpretation of $\beta$: it is inversely proportional to the (unbiased) transition rates between minima of the potential. [^3] From the Kramer formula; $\text{transition rate}\propto\frac{1}{\beta}\exp(\beta\Delta V)$, where $\Delta V\propto\frac{1}{\beta^{\prime}}=\frac{1}{\beta}$. We also set the kernel width equal to the width of the Gaussian bias. This is reasonable since it is the variability of $V(x)$ that determines sensible values for both these parameters. The resulting exploration algorithm we call Adapted Metadynamics (AM), and is presented in Algorithm 1. Note that the algorithm can be extended to a batch of trajectories, where each metadynamics trajectory evolves independently, but with a shared $\hat{V}$ and $V_{\text{bias}}$ which receive updates from every trajectory in the batch. This is the version we use in our experiments - it accelerates exploration and reduces stochastic gradient noise during training.
Input : Manifold environment of terminating states $\mathcal{X}$ with reward density $r\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. Initial state $(x_{t},p_{t})\in\mathcal{X}\times T_{x_{t}}(\mathcal{X})$. Collective variables $z=\left(z_{1},\dots,z_{k}\right)$.
Parameters : Gaussian width $\sigma=\left(\sigma_{1},\dots,\sigma_{k}\right)\in\mathbb{R}^{k}$. Gaussian height $w>0$. Stride $n\in\mathbb{Z}^{+}$. LD parameters: $\gamma,\beta$. Timestep $\Delta t$.
1
[…]
3 $\hat{R}\leftarrow 0$
4 $\hat{V}(z)\leftarrow 0$
5 $V_{\text{bias}}(z)\leftarrow 0$
6 every timestep $\Delta t$:
7   $z_{t}\leftarrow z(x_{t})$
8   every n timesteps $n\Delta t$:
9     $\hat{N}\leftarrow\hat{N}+\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
10     $\hat{R}\leftarrow\hat{R}+r(x_{t})\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
11     $\hat{V}\leftarrow-\frac{1}{\beta}\log\left(\hat{R}/(\hat{N}+\epsilon)+\epsilon\right)$
12     $V_{\text{bias}}(z)\leftarrow V_{\text{bias}}(z)+n\cdot\Delta t\cdot w\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
13   compute forces:
14     $F\leftarrow-\left(\nabla_{z}\hat{V}(z)|_{z=z_{t}}+\nabla_{z}V_{\text{bias}}(z)|_{z=z_{t}}\right)\cdot\nabla_{x}z|_{x=x_{t}}$
  propagate $x_{t},p_{t}$ by $\Delta t$ using Langevin dynamics with computed force $F$ (Alg. 2, Appendix D).
Algorithm 1 Adapted Metadynamics
Training GFlowNets with Adaptive Metadynamics (MetaGFN) Each Adaptive Metadynamics sample $x_{i}\in\mathcal{X}$ is an off-policy terminal state sample. To train a GFlowNet, complete trajectories are required. We generate these by backward sampling from the terminal state, giving a trajectory $\tau=(s_{0},s_{1},\ldots,s_{n}=x_{i})$, where each state $s_{i-1}$ is sampled from the current backward policy distribution $\hat{p}_{B}(s_{i-1}|s_{i};\theta)$, for $i$ from $n$ to $1$. This approach means that the generated trajectory $\tau$ has reasonable credit according to the loss function, thereby providing a useful learning signal. However, since this requires a backward policy, this is compatible with DB, TB, and STB losses, but not FM loss. Given the superior credit assignment of the former losses, this is not a limitation Madan et al. (2022).
Additionally, we use a replay buffer. Due to the theoretical guarantee that Adaptive Metadynamics will eventually sample all collective variable space (Theorem 3.1), AM samples are ideal candidates for storing in a replay buffer. When storing these trajectories in the replay buffer, there are two obvious choices:
  1. 1.
    Store the entire trajectory the first time it is generated;
  2. 2.
    Store only the Adaptive Metadynamics sample and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
We investigate both options in our experiments. We call the overall training algorithm MetaGFN, with pseudocode presented in Algorithm 3, Appendix D.
4 Experiments
4.1 Line environment
Figure 1: Line Environment reward function, equation (2).
We consider a one-dimensional line environment, with state space $\mathcal{S}=\mathbb{R}\times\left\{t\in\mathbb{N},1\leq t\leq 3\right\}$, where $t$ indexes the position of a state in a trajectory. The source state is $s_{0}=(0,0)$ and trajectories terminate after exactly $3$ steps. The terminal states are therefore $\mathcal{X}=\mathbb{R}\times\left\{3\right\}\cong\mathbb{R}$. The reward density, plotted in Figure 1, consists of an asymmetric bimodal peak near the origin and an additional distant lone peak. It is given by the Gaussian mixture distribution:
$r(x)=\begin{cases}\mathcal{N}(-2.0,1.0)+\mathcal{N}(-2.0,0.4)+\\ \quad\mathcal{N}(2.0,0.6)+\mathcal{N}(20.0,0.1);&-5\leq x\leq 23\\ 0;&\text{otherwise},\end{cases}$ (2)
where $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian density with mean $\mu$ and variance $\sigma^{2}$. The forward and backward probability transition kernels are a mixture of $3$ Gaussian distributions which, along with the flow $\hat{f}$, are parameterised by an MLP.
We compare the following exploration strategies: entirely on-policy (no exploration), noisy exploration, Thompson Sampling, and MetaGFN. For each strategy, we use a replay buffer and alternate between an exploration batch and a replay buffer batch. For MetaGFN, we use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$ and always regenerate MetaGFN training trajectories using the current backward policy. As the environment is one-dimensional, the collective variable is simply $z(x)=x$. We evaluate performance by computing the L1 error between the known reward distribution and the empirical on-policy distribution during training (see Appendix E.1 for the full experimental details).
Figure 2: The L1 difference between on-policy and reward distribution during training in the line environment for different loss functions and exploration strategies. The mean is plotted with standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 2. Among the three loss functions, TB loss has the lowest variance loss profiles, and MetaGFN consistently converges to a lower minimum error than all other exploration strategies. Indeed, MetaGFN was the only method that consistently sampled the distant reward peak at $x=20$, while other methods plateau in error after locking onto the central modes (Appendix E.2). Despite this, we observed that in the initial stages of training, all exploration strategies found occasional samples from the distant peak. The reason MetaGFN is the only method that converges is that Adaptive Metadynamics manages to consistently sample the distant peak, even if the on-policy starts to focus on the central modes. This keeps the replay buffer populated with samples from every reward peak during training, which eventually encourages the on-policy to sample from every mode. The small increase in the loss of MetaGFN around batch number $5\times 10^{3}$ happens because the on-policy distribution widens when Adaptive Metadynamics first discovers the distant peak. In Appendix E.2, we show further details of Adaptive Metadynamics in this environment and we compare different MetaGFN variants, with and without noise, and with and without trajectory regeneration. We confirm that the version of MetaGFN presented in Figure 2 (no added noise and always regenerate trajectories) is the most robust variant.
4.2 Grid environment
Figure 3: Grid Environment reward function.
We consider a two-dimensional environment, with state space $\mathcal{S}=\mathbb{R}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$. The source state is $s_{0}=(0,0,0)$ and trajectories terminate after exactly $3$ steps. Terminal states are $\mathcal{X}=\mathbb{R}^{2}\times\{3\}\cong\mathbb{R}^{2}$. The reward function is given by the sum of four symmetric Gaussians with means $\mu=(-7,-7),(-7,7),(7,7),(7,-7)$ and variances $\sigma^{2}=2$ (Figure 3). The forward and backward probability transition kernels are a mixture of $4$ bivariate Gaussian distributions with means and diagonal variance matrices, along with the flow $\hat{f}$, parameterised by an MLP. We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix F for full experiment details).
Figure 4: The L1 difference between on-policy and reward distribution during training in the grid environment for different loss functions and exploration strategies. The mean is plotted with standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 4. We see that for most loss functions and exploration strategies (except Thompson sampling with TB loss), the training curves converge to a low loss, indicating that all four modes of the grid environment are well-sampled. However, in all cases, MetaGFN converges faster than other exploration strategies. The fastest convergence occurs for MetaGFN with TB loss; after $10^{3}$ batches MetaGFN has reached a similar loss that noisy exploration reaches after $10^{4}$ batches and on-policy reaches after $\sim 3\times 10^{4}$ batches. Meanwhile, the loss curves have lower error indicating more stable training with MetaGFN in this environment.
4.3 Alanine dipeptide environment
Figure 5: (a) Free energy surface of alanine dipeptide in an explicit solvent with all metastable states annotated. (b) The potential KDE $\hat{V}$ learnt using Adapted Metadynamics with von Mises kernel, concentration $\kappa=10$.
One application of continuous GFlowNets is molecular conformation sampling Volokhova et al. (2023). Here, we train a GFlowNet to sample conformational states of alanine dipeptide (AD), a small biomolecule of $23$ atoms that plays a key role in modelling backbone dynamics of proteins Hermans (2011). The metastable states of AD can be distinguished in a two-dimensional CV space of $\phi$ and $\psi$, the two backbone dihedral angles. The resulting free energy surface for AD in explicit water, $V(\phi,\psi)$, obtained after extensive sampling long molecular dynamics simulation is shown in Figure 5(a). The metastable states (energy minima), in increasing energy, are $P_{||}$, $\alpha_{R}$, $C_{5}$, $\alpha^{\prime}$, $\alpha_{L}$ and $\alpha_{D}$.
The state space is $\mathcal{S}=\mathbb{T}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$. The source state is $s_{0}=P_{||}=(-1.2,2.68)$ and trajectories terminate after exactly $3$ steps. Terminal states are $\mathcal{X}=\mathbb{T}^{2}\times\{3\}\cong\mathbb{T}^{2}$. The reward density is given by the Boltzmann weight, $r(\phi,\psi)=\frac{1}{Z}\exp(-\beta V(\phi,\psi))$, where $Z$ is the normalisation constant. The forward and backward probability transition kernels are defined as a mixture of $3$ bivariate von Mises distributions, parameterised through an MLP. We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix G for full experimental details).
Figure 6: The L1 difference between on-policy and reward distribution during training in the alanine dipeptide environment for different loss functions and exploration strategies. The mean is plotted with a standard error over 10 repeats. DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
Figure 7: Learned on-policy distribution for TB on-policy and TB MetaGFN training runs. The colour bar shows the probability density. Red histograms show the marginal distribution along the angular coordinates. Black curves show the marginal distributions of the ground truth. In the best case, MetaGFN is able to learn the $\alpha_{L}$ mode. In the worst case, MetaGFN fails to converge. On-policy training, although more consistent, fails to learn to sample from the $\alpha_{L}$ mode.
The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 6. For each loss function, models trained with MetaGFN generally converge to a lower minimum error than all other exploration strategies. For TB loss, however, the average L1 error is marginally higher than on-policy training, but this conceals the fact that the best-case error is smaller. Thus, to better understand this result, we examine the best and worst training runs (as measured by L1 error) for TB on-policy and TB MetaGFN, shown in Figure 7. We see that the best run trained with metadynamics can sample from the rare $\alpha_{L}$ minima, unlike the on-policy run. In the worst run, MetaGFN fails to converge (although this is rare; only one of 10 runs failed). In Table 1, we quantify how often the different AD modes were correctly sampled over the different repeats (a mode is correctly sampled if the on-policy distribution also has a mode within the correct basin of attraction). The only mode not correctly sampled by any method is $\alpha_{D}$, which has a natural abundance approximately $10$ times less frequent than $\alpha_{L}$. We see that TB loss with MetaGFN is the only combination that can consistently sample the majority of modes, whilst noisy exploration and Thompson Sampling both perform worse than on-policy in this environment.
Table 1: Number of correct samples of AD modes in trained GFlowNets over 10 independent repeats for DB, STB, and TB loss functions. OP - On-policy and MD - MetaGFN. The $\alpha_{D}$ mode wasn’t sampled in any model due to its low natural frequency.
DB STB TB
OP MD OP MD OP MD
$P_{||}$ 1 7 6 5 10 8
$\alpha_{R}$ 6 9 7 10 10 9
$C_{5}$ 2 7 5 6 10 8
$\alpha^{\prime}$ 6 9 5 10 5 9
$\alpha_{L}$ 0 1 1 0 0 8
5 Limitations
For metadynamics to be an effective sampler, the CVs must be low-dimensional and bounded, properties that were satisfied in both our experimental environments. Therefore, it is necessary to either know such CVs in advance, assuming they exist or learn them automatically from data Sidky et al. (2020). An alternative approach would be to learn CVs adaptively by parameterising the CV function by a neural network and updating its parameters by back-propagating through the GFlowNet loss when training on MetaGFN trajectories. A final improvement could be to replace the metadynamics algorithm itself with a variant with smoother convergence properties, such as well-tempered metadynamics Barducci et al. (2008) or on-the-fly probability enhanced sampling (OPES) Invernizzi (2021). We leave these extensions for future work.
6 Conclusions
While exploration strategies for discrete Generative Flow Networks (GFlowNets) have received extensive attention, the methodologies for continuous GFlowNets remain relatively underexplored. To address this gap, we illustrated how metadynamics, a widely used enhanced sampling technique in molecular dynamics, can be adapted as an effective exploration strategy for continuous GFlowNets.
In molecular dynamics, atomic forces can be computed as the gradient of the potential, whereas continuous GFlowNets tackle problems where the reward function is a black box and gradients are inaccessible. We demonstrated how the method could be adapted by updating a kernel density estimate of the reward function on-the-fly, and proved that this is guaranteed to explore the space in an appropriate limit. Our empirical investigations show that MetaGFN offers a computationally efficient means to explore new modes in environments where prior knowledge of collective variables exists. Importantly, this work advocates an approach wherein techniques derived from molecular modelling can be adapted for machine learning tasks. Looking ahead, we anticipate that this could be a fruitful area of cross-disciplinary research, where existing ideas from the enhanced sampling literature can find applications in a broad range of generative modelling and reinforcement learning tasks.
Appendix A Loss functions
For a complete trajectory $\tau$, the detailed balanced loss (DB) is
$L_{DB}(\tau)=\sum_{t=0}^{n-1}\left(\log\frac{\hat{f}(s_{t};\theta)\hat{p}_{F}(s_{t},s_{t+1};\theta)}{\hat{f}(s_{t+1};\theta)\hat{p}_{B}(s_{t+1},s_{t};\theta)}\right)^{2},$
where $\hat{f}(s_{t+1};\theta)$ is replaced with $r(s_{n})$ if $s_{n}$ is terminal.
The subtrajectory balance loss (STB) is
$L_{STB}(\tau)=\frac{\sum_{0\leq i<j\leq n}\lambda^{j-1}\mathcal{L}_{TB}(\tau_{i:j})}{\sum_{0\leq i<j\leq n}\lambda^{j-i}},$
$\mathcal{L}_{STB}(\tau_{i:j})\vcentcolon=\left(\log\frac{\hat{f}(s_{i};\theta)\prod_{t=i}^{j-1}\hat{p}_{F}(s_{t+1}|s_{t};\theta)}{\hat{f}(s_{j};\theta)\prod_{t=i+1}^{j}\hat{p}_{B}(s_{t-1}|s_{t};\theta)}\right)^{2},$
where $\hat{f}(s_{j};\theta)$ is replaced with $r(s_{j})$ if $s_{j}$ is terminal. In the above, $\lambda<0$ is a hyperparameter. The limit $\lambda\rightarrow 0^{+}$ leads to average detailed balance. The $\lambda\rightarrow\infty$ limit gives the trajectory balance objective. We use $\lambda=0.9$ in our experiments.
Appendix B Langevin dynamics
Langevin dynamics (LD), is defined through the Stochastic Differential Equation (SDE):
$\mathrm{d}x=M^{-1}p\mathrm{d}t$ (3)
$\mathrm{d}p=F(x)\mathrm{d}t-\gamma p\mathrm{d}t+\sqrt{2\gamma\beta^{-1}}M^{1/2}\mathrm{d}W.$ (4)
In the above, $x,p\in\mathbb{R}^{D}$ are vectors of instantaneous position and momenta respectively, $F\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ is a force function, $W(t)$ is a vector of $D$ independent Wiener processes, $M\in\mathbb{R}^{D}\times\mathbb{R}^{D}$ is a constant diagonal mass matrix, and $\gamma,\beta>0$ are constant scalars which can be interpreted as a friction coefficient and inverse temperature respectively. In conventional Langevin dynamics, the force function is given by the gradient of the potential energy function, $F=-\nabla V$, where $V\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the dynamics are ergodic with respect to the Gibbs-Boltzmann density
$\rho_{\beta}(x,p)\propto e^{-\beta H(x,p)},$
where $H(x,p)=p^{T}M^{-1}p/2+V(x)$ is the Hamiltonian. Since the Hamiltonian is separable in position and momenta terms, the marginal Gibbs-Boltzmann density is position space is simply $\rho_{\beta}(x)\propto e^{-\beta V(x)}$.
Appendix C Proofs
Lemma C.1.
Let $(f_{n}(x))$ and $(g_{n}(x))$ be sequences of real functions where $\lim_{n\rightarrow\infty}f_{n}(x)=\infty$, $\lim_{n\rightarrow\infty}g_{n}(x)=\infty$ and $\lim_{n\rightarrow\infty}\frac{f_{n}(x)}{g_{n}(x)}=h(x)$. Then, for all $\epsilon>0$, we have $\lim_{n\rightarrow\infty}\frac{f_{n}(x)}{g_{n}(x)+\epsilon}=h(x)$.
Proof.
$\lim_{n\rightarrow\infty}\frac{f_{n}(x)}{g_{n}(x)+\epsilon}=\lim_{n\rightarrow\infty}\frac{f_{n}(x)}{g_{n}(x)}\frac{1}{1+\epsilon/g_{n}(x)}$
and the right-hand side is the product of two functions whose limit exists so, by the product rule of limits
$\lim_{n\rightarrow\infty}\frac{f_{n}(x)}{g_{n}(x)}\lim_{n\rightarrow\infty}\frac{1}{1+\epsilon/g_{n}(x)}=h(x)\cdot 1=h(x),$
so done. ∎
Lemma C.2.
Let $(X_{i})$ be a sequence of continuous random variables that take values on a bounded domain $D\subset\mathbb{R}^{d}$ that asymptotically approaches the uniform random variable $U$ on $D$, i.e. $X_{i}\rightarrow U$ uniformly. Further, suppose
$h(x)\vcentcolon=\frac{\sum_{i=1}^{\infty}f(x_{i})g(x,x_{i})}{\sum_{i=1}^{\infty}g(x,x_{i})}$
exists, where $x_{i}\in D$ is a sample from $X_{i}$ and $f(x)$ and $g(x,x^{\prime})$ are analytic functions on $D$ and $D\times D$ respectively. Then,
$h(x)=\frac{\sum_{i=1}^{\infty}f(u_{i})g(x,u_{i})}{\sum_{i=1}^{\infty}g(x,u_{i})},$
where the $u_{i}$ are samples from $U$. We make no assumption of independence of samples.
Proof.
Fix a probability space $(\Omega,\mathcal{F},P)$ on which $(X_{i})$ and $U$ are defined. Recall that a continuous random variable $X$ that takes values on $D\subset\mathbb{R}^{d}$ is a measurable function $X:\Omega\rightarrow D$ where $(D,\mathcal{B})$ is a measure space and $\mathcal{B}$ is the Borel $\sigma$-algebra on $D$. Let $\omega\subset\Omega$ denote an arbitrary element of the sample space. The requirement that $X_{i}\rightarrow U$ uniformly can be written formally as:
$\forall\epsilon>0,\exists N(\epsilon)\text{ s.t. }\forall i>N(\epsilon),\forall\omega\subset\Omega,|X_{i}(\omega)-U(\omega)|<\epsilon.$
We prove the Lemma by showing that equality holds for all possible sequences of outcomes $\omega_{1},\omega_{2},\dots$. That is, we prove:
$\frac{\sum_{i=1}^{\infty}f(X_{i}(\omega_{i}))g(x,X_{i}(\omega_{i}))}{\sum_{i=1}^{\infty}g(x,X_{i}(\omega_{i}))}=\frac{\sum_{i=1}^{\infty}f(U(\omega_{i}))g(x,U(\omega_{i})))}{\sum_{i=1}^{\infty}g(x,U(\omega_{i})))}.$ (5)
Since these are ratios of infinite series, to prove their equality it is sufficient to show that the numerator of the LHS is asymptotically equivalent to the numerator of the RHS and that the denominator of the LHS is asymptotically equivalent to the denominator of the RHS. Recall that two sequences of real functions $(a_{n})$ and $(b_{n})$ are asymptotically equivalent if $\lim_{n\rightarrow\infty}\frac{a_{n}(x)}{b_{n}(x)}=c$ where $c$ is a constant. First, we prove that this holds with
$a_{n}\vcentcolon=\sum_{i=1}^{n}f(X_{i}(\omega_{i}))g(x,X_{i}(\omega_{i}))$ (6)
and
$b_{n}\vcentcolon=\sum_{i=1}^{n}f(U(\omega_{i}))g(x,U(\omega_{i}))).$ (7)
We write
$\frac{a_{n}}{b_{n}}=\frac{\sum_{i=1}^{N(\epsilon)}f(X_{i}(\omega_{i}))g(x,X_{i}(\omega_{i}))+\sum_{i=N(\epsilon)+1}^{n}f(X_{i}(\omega_{i}))g(x,X_{i}(\omega_{i}))}{\sum_{i=1}^{N(\epsilon)}f(U(\omega_{i}))g(x,U(\omega_{i})))+\sum_{i=N(\epsilon)+1}^{n}f(U(\omega_{i}))g(x,U(\omega_{i})))}.$
Dividing by $\sum_{i=N(\epsilon)+1}^{n}f(U(\omega_{i}))g(x,U(\omega_{i})))$ and taking the limit $n\rightarrow\infty$ we have
$\lim_{n\rightarrow\infty}\frac{a_{n}}{b_{n}}=\lim_{n\rightarrow\infty}\frac{\sum_{i=N(\epsilon)+1}^{n}f(X_{i}(\omega_{i}))g(x,X_{i}(\omega_{i}))}{\sum_{i=N(\epsilon)+1}^{n}f(U(\omega_{i}))g(x,U(\omega_{i})))}.$
Since $f$ and $g$ are analytic and $i>N(\epsilon)$ for all terms in the sums we have, by Taylor expansion, $f(X_{i}(\omega_{i}))=f(U(\omega_{i}))+O(\epsilon)$ and $g(x,X_{i}(\omega_{i}))=g(x,U(\omega_{i}))+O(\epsilon)$, hence
$\lim_{n\rightarrow\infty}\frac{a_{n}}{b_{n}}=\lim_{n\rightarrow\infty}\left(1+\frac{nO(\epsilon)}{\sum_{i=N(\epsilon)+1}^{n}f(U(\omega_{i}))g(x,U(\omega_{i})))}\right)=1+\lim_{n\rightarrow\infty}\frac{nO(\epsilon)}{O(n)}=1+O(\epsilon).$
Finally, since $\epsilon$ can be made arbitrarily small by partitioning the sum at a $N(\epsilon)$ that is sufficiently large, we conclude that $\lim_{n\rightarrow\infty}\frac{a_{n}}{b_{n}}=1$, hence $(a_{n})$ and $(b_{n})$ as defined in (6) and (7) are asymptotically equivalent. By a similar argument, it can be shown that
$c_{n}\vcentcolon=\sum_{i=1}^{n}g(x,X_{i}(\omega_{i}))$
and
$d_{n}\vcentcolon=\sum_{i=1}^{n}g(x,U(\omega_{i})))$
are also asymptotically equivalent. This proves (5).
Below, we present the proof of Theorem 3.1 that appears in the main text.
Proof.
For concreteness, throughout this proof we assume that the kernel function is a Gaussian. We explain at the appropriate stage in the proof, indicated by (*), how this assumption can be relaxed.
First we take the $t\rightarrow\infty$ limit. Since the log function is continuous, the limit and log can be interchanged and we have
$\lim_{t\rightarrow\infty}\hat{V}(x,t)=-\frac{1}{\beta^{\prime}}\log\left(\lim_{t\rightarrow\infty}\left(\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}\right)+\epsilon\right).$
Recall the uniform time discretisation $t_{n}=n\Delta t$ of metadynamics (Algorithm 1). Thus, we can write
$\hat{R}(x,t_{n})=\sum_{i=1}^{n}r(x_{i})\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)},$ (8)
$\hat{N}(x,t_{n})=\sum_{i=1}^{n}\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$ (9)
Since the domain is bounded, we know that for fixed $x$, both (8) and (9) have limit at infinity, i.e. $\lim_{t\rightarrow\infty}\hat{R}(x,t)=\infty$ and $\lim_{t\rightarrow\infty}\hat{N}(x,t)=\infty$. Hence, by Lemma C.1, we have
$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}=\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)},$
provided the limit on the RHS exists. Next, we show that this limit exists by computing it explicitly. The limit can be written
$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)}=\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(x_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
Recall that metadynamics eventually leads to uniform sampling over the domain, independent of the potential. Hence, since $R$ and $z(x)$ are analytic, by Lemma C.2 we may replace the metadynamics samples $x_{i}$ with samples from a uniform distribution, denoted $u_{i}$:
$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(u_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
In the limit, the ratio of sums with uniform sampling becomes a ratio of integrals:
$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\frac{\int_{D}r(x^{\prime})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}}{\int_{D}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}},$
where $D\subset\mathbb{R}^{d}$ is the domain. The limit is therefore a (scaled) convolution of the reward function with a Gaussian in the collective variable space with width vector $\sigma$. Taking the limit $\sigma_{i}\rightarrow 0$ for all $i\in\left\{1,2,\dots,d\right\}$, the Gaussian convergences to a delta distribution in the collective variable space and we have
$\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}.$
(*) This step also holds for any kernel that becomes distributionally equivalent to a Dirac delta function in the limit that its variance parameter goes to zero. In particular, it also holds for the von Mises distribution that we use in our alanine dipeptide experiment in $\mathbb{T}^{2}$.
Finally, we take the limit $\epsilon\rightarrow 0$ to obtain
$\lim_{\epsilon\rightarrow 0}\lim_{\sigma\rightarrow 0}\lim_{t\rightarrow\infty}\hat{V}(x,t)$ $=-\frac{1}{\beta^{\prime}}\lim_{\epsilon\rightarrow 0}\log\left(\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}+\epsilon\right)$ (10)
$=-\frac{1}{\beta^{\prime}}\int_{D}\log\left(r(x^{\prime})\right)\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}$ (11)
$=\int_{D}V(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}\vcentcolon=V(z),$ (12)
where we have used the definition $V(x^{\prime})=-\frac{1}{\beta^{\prime}}\log(r(x^{\prime}))$ and in the last step we used the definition of the marginal potential energy in the collective variable space. If $z(x)$ is invertible, then the delta function simplifies to a delta function in the original space and we obtain the original potential instead of the marginal potential. ∎
Appendix D Algorithms
Input : Current state $(x_{t},p_{t})$. Force $F$.
Parameters : Friction coefficient $\gamma$. Thermodynamic beta $\beta$. Timestep $\Delta t$.
Output : State $(x_{t+\Delta t},p_{t+\Delta t})$ at the next timestep.
1
2Sample a random vector $R$, with the same dimension as $x_{t}$, where each element is an independent sample from a standard normal.
3 $x_{t+\Delta t}=x_{t}+p_{t}\Delta t$
4 $p_{t+\Delta t}=p_{t}+F\Delta t-\gamma p_{t}\Delta t+\sqrt{2\gamma\Delta t/\beta}\cdot R$
5return $(x_{t+\Delta t},p_{t+\Delta t})$
Algorithm 2 Euler-Maruyama Langevin Dynamics Step
In the algorithm below, we present the variant of MetaGFN where we store Adaptive Metadynamics samples in the replay buffer and regenerate trajectories using the current backward policy when retrieving from the replay buffer. This is the variant we used in our experiments in the main text.
Input : Forward policy $P_{F}$. Backwards policy $P_{B}$. Loss function $L$.
Parameters : How often to run Adaptive Metadynamics batches, freqMD. How often to run replay buffer batches, freqRB. Batch size, $b$. Stride, $n\in\mathbb{Z}^{+}$. Time step, $\Delta t>0$
1
2for each episode do:
3if episode number is divisible by freqMD:
4   Run Adaptive Metadynamics (batch size $b$) for time $n\Delta t$, obtain samples $\{x_{1},\dots,x_{b}\}$
5   Push $\{x_{1},\dots,x_{b}\}$ to the replay buffer
6   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
7elif episode number is divisible by freqRB:
8   Random sample $\{x_{1},\dots,x_{b}\}$ from the replay buffer
9   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
10
11else:
12   Generate trajectories $\{\tau_{1},\dots,\tau_{b}\}$ on-policy
13
14 Compute loss $l=\sum_{i=1}^{b}L(\tau_{i},P_{F},P_{B})$
 Take gradient step on loss $l$
Algorithm 3 MetaGFN
Appendix E Experiment details: line environment
E.1 Experimental setup
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 256 hidden units per layer. We use the GELU activation function and dropout probability $0.2$ after each layer. This defines the torso of the MLP. Connecting from this common torso, the MLP has three single-layer, fully connected heads. The first two heads have output dimension $9$ and parameterise the $3$ means ($\mu$), standard deviations ($\sigma$) and weights ($w$) of the mixture of Gaussians for the forward and backward policies respectively. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The mean and standard deviation outputs are passed through a sigmoid function and transformed so that they map to the ranges $\mu\in(-14,14)$ and $\sigma\in(0.1,1)$. The mixture weights are normalised with the softmax function. The exception to this parameterisation is the backward transition to the source state which is fixed to be the Dirac delta distribution centred on the source, i.e. $\hat{p}_{B}(s_{0}|s_{1};\theta)=\delta_{s_{0}}$. For the TB loss, we treat $\log Z_{\theta}$ as a separate learnable parameter.
Replay buffer The replay buffer has capacity for $10^{4}$ trajectories. Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-3}$. When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
Figure 8: Exponential noise schedule.
Noisy exploration Noisy exploration is defined by adding an additional constant, $\bar{\sigma}$, to the standard deviations of the Gaussian distributions of the forward and backward policies. Specifically, the forward policy becomes $\hat{p}_{F}(s_{t},s_{t-1};\theta)=\sum_{i=1}^{3}w_{i}\mathcal{N}(\mu_{i},(\sigma_{i}+\bar{\sigma})^{2})$, and similarly for the backward policy. We schedule the value of $\bar{\sigma}$ so that it decreases during training according to an exponential-flat schedule:
$\bar{\sigma}=\begin{cases}\bar{\sigma}_{0}\left(e^{-2je/(B/2)}-e^{-2e}\right)\;\;&j<B/2\\ 0\;\;&j\geq B/2,\end{cases}$ (13)
where $j\in(1,\dots,B)$ is the batch number and $\bar{\sigma}_{0}=2$ is the initial noise, plotted in Figure 8.
Thompson Sampling We use $10$ heads with the bootstrapping probability parameter set to $0.3$.
MetaGFN We use $\Delta t=0.05$, $n=2$, $\beta=1$, $\gamma=2$, $w=0.15$, $\sigma=0.1$, $\epsilon=10^{-3}$. The domain of Adaptive Metadynamics is restricted to $[-5,23]$ and reflection conditions are imposed at the boundary. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.01$. Initial metadynamics samples are drawn from a Gaussian distribution, mean 00 and variance $1$, and initial momenta from a Gaussian distribution, mean 00 and variance $0.5$.
Training parameters In all experiments, we use batch size $b=64$ and for $B=10^{5}$ batches. We use a learning rate with a linear schedule, starting at $10^{-3}$ and finishing at 00. For the TB loss, we train the $\log Z_{\theta}$ with a higher initial learning rate of $10^{-1}$ (also linearly scheduled). For the STB loss, we use $\lambda=0.9$, a value that has worked well in the discrete setting Madan et al. [2022]. We use the Adam optimiser with gradient clipping. For all loss functions, we clip the minimum log-reward signal at $-10$. This enables the model to learn despite regions of near-zero reward between the modes of $r(x)$.
Evaluation The L1 error between the known reward distribution, $\rho(x)=r(x)/Z$, and the empirical on-policy distribution, denoted $\hat{\rho}(x)$, estimated by sampling $10^{4}$ on-policy trajectories and computing the empirical distribution over terminal states. Specifically, we compute
$\text{error}=\frac{1}{2}\int_{-5}^{23}\left|\hat{\rho}(x)-\frac{r(x)}{Z}\right|\mathrm{d}x,$ (14)
where the integral is estimated by a discrete sum with grid spacing $0.01$. Note that this error is normalised such that for all valid probability distributions $\hat{\rho}(x)$, we have $0\leq\text{error}\leq 1$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $1$ hour to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics was negligible compared to the training time of the models.
Figure 9: Forward policy and replay buffer distributions after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can consistently learn all three peaks.
E.2 Results
On-policy distributions Figure 9 shows the forward policy and replay buffer distributions (with bias sampling) after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can uniformly populate the replay buffer and consistently learn all three peaks.
Adaptive Metadynamics Figure 10(a) shows the L1 error between the density implied by the KDE potential and the true reward distribution during a typical training run. Figure 10(b) shows the resulting $\hat{V}$ and $V_{\text{bias}}$ at the end of the training. By (1), Adaptive Metadynamics has fully explored the central peaks. At (2), the third peak is discovered, prompting rapid adjustment of the KDE potential. By $2.5\times 10^{4}$ iterations, a steady state is reached and the algorithm is sampling the domain uniformly.
Figure 10: (a) L1 error between $\hat{\rho}=\exp(-\beta\hat{V}(x))/Z$ and the reward distribution, $r(x)/Z$ (b) kde potential, bias potential, and positions of final samples after $2.5\times 10^{4}$ training iterations.
Comparing MetaGFN Variants We consider three MetaGFN variants. The first variant, always backwards sample, regenerates the entire trajectory using the current backward policy when pulling from the replay buffer. The second variant, reuse initial backwards sample, generates the trajectory when first added to the replay buffer and reuses the entire trajectory if subsequently sampled. The third variant, with noise, is to always backwards sample with noisy exploration as per equation (13). We plot the L1 policy errors in Figure 11. We observe that always backward sample is better than reuse initial backwards sample for all loss functions. For DB and TB losses, there is no evidence of any benefit of adding noise, whereas noise improves training for STB loss, performing very similarly to TB loss without noise.
Figure 11: Comparing MetaGFN variants.
Appendix F Experiment details: grid environment
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 512 hidden units per layer, GELU activation function and dropout probability $0.2$, similar to the Line Environment. The first two heads have output dimension $24$ and parameterise the $8$ means, $12$ standard deviations and $4$ weights of the mixture of bivariate Gaussians. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The means and standard deviations are mapped to the range $\mu\in(-15,15)$ and $\sigma\in(0.1,7)$ through sigmoid functions. Mixture weights are normalised with the softmax function.
Replay buffer Identical to the Line Environment replay buffer, except trajectories are only stored when the terminal state’s reward exceeds $10^{-4}$.
Noisy exploration We use the same noise profile as the Line Environment, equation (13). The noise $\bar{\sigma}$ is added equally to both dimensions of the standard deviation.
Thompson sampling We use $10$ heads with the bootstrap probability parameter set to $0.3$.
MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.35$, $n=3$, $\beta=1$, $\gamma=2$, $w=0.10$, $\sigma=2$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.075$. Initial samples are drawn from a Gaussian distribution, mean centred $(0,0)$, variance $\sigma^{2}=(1,1)$. Momenta are initialised to zero.
Training parameters The same as for the Line Environment, see Appendix E.1.
Evaluation The L1 error between the histogram of the terminal states of $10^{4}$ on-policy samples and the exact distribution is computed. $\text{error}=\frac{1}{2}\int_{-15}^{15}\int_{-15}^{15}\left|\hat{\rho}(x,y)-\frac{r(x,y)}{Z}\right|\mathrm{d}x\mathrm{d}y$, estimated by a discrete sum with grid spacing $0.075$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics is less than $10\%$.
Appendix G Experiment details: alanine dipeptide environment
Computing the Free Energy Surface To obtain a ground-truth free energy surface (FES) in $\phi$-$\psi$ space, we ran a 250ns NPT well-tempered metadynamics MD simulation of alanine dipeptide at temperature 300K ($\beta=0.4009$), pressure 1bar with the TIP3P explicit water model Jorgensen et al. [1983]. We used the PLUMMED plugin Bonomi et al. [2019] for OpenMM Eastman et al. [2017] with the AMBER14 force field Salomon-Ferrer et al. [2013].
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and $\hat{f}$ through three heads of an MLP with 3 hidden layers with 512 hidden units per layer, with GeLU activations and dropout probability $0.2$, similar to the Line Environment. The first two heads have output dimensions of $15$, parameterising the $6$ means, $6$ concentrations, and $3$ weights of the mixture of von Mises policy. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The means are mapped to the range $(-\pi,\pi)$ through $2\arctan(\cdot)$. Concentrations are parameterised in log space and passed through a sigmoid to map to the range $\ln(\kappa)\in(0,5)$. Mixture weights are normalised with the softmax function.
Replay buffer The replay buffer has capacity for $10^{4}$ trajectories. Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-10}$. When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
Noisy exploration We use the same noise profile as the Line Environment, equation (13). The noise $\bar{\sigma}$ is now added to the concentration parameter $\kappa$. Concentration is related to standard deviation through $\sigma=\frac{1}{\kappa^{2}}$.
Thompson Sampling We use 10 heads with the bootstrap probability parameter set to 0.3.
MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.01$, $n=2$, $\beta=0.4009$, $\gamma=0.1$, $w=10^{-5}$, $\kappa=10$, $\epsilon=10^{-6}$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.1$. Initial samples are drawn from a Gaussian distribution, mean centred $P_{||}$, variance $\sigma^{2}=(0.1,0.1)$, and initial momenta from a Gaussian, mean $\mu=(0,0)$, variance $\sigma^{2}=(0.05,0.05)$.
Training parameters The same as for the Line Environment, see Appendix E.1.
Evaluation The L1 error of a histogram of on-policy samples, $\hat{\rho}(\phi,\psi)$, is computed via a two-dimensional generalisation of (14); $\text{error}=\frac{1}{2}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left|\hat{\rho}(\phi,\psi)-\frac{r(\phi,\psi)}{Z}\right|\mathrm{d}\phi\mathrm{d}\psi$, estimated by a discrete sum with grid spacing $0.1$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics was less than $10\%$.

### v2

@@
 We use Adapted Metadynamics as an exploration strategy for continuous GFlowNets.
-We show three continuous domains where the resulting algorithm, MetaGFN, accelerates convergence to the target distribution and discovers more distant reward modes than previous off-policy exploration strategies used for GFlowNets.
+We show several continuous domains where the resulting algorithm, MetaGFN, accelerates convergence to the target distribution and discovers more distant reward modes than previous off-policy exploration strategies used for GFlowNets.
 1 Introduction
-Generative Flow Networks (GFlowNets) are a type of generative model that samples from a discrete space $\mathcal{X}$ by sequentially constructing objects via actions taken from a learned policy $P_{F}$ Bengio et al.
-(2021a).
+Generative Flow Networks (GFlowNets) are a type of generative model that samples from a discrete space $\mathcal{X}$ by sequentially constructing objects via actions taken from a learned policy $P_{F}$ [1].
 The policy $P_{F}(s,s^{\prime})$ specifies the probability of transitioning from some state $s$ to some other state $s^{\prime}$.
@@
 GFlowNets offer advantages over more traditional sampling methods, such as Markov chain Monte Carlo (MCMC), by learning an amortised sampler, capable of single-shot generation of samples from the desired distribution.
-Since GFlowNets learn a parametric policy, they can generalise across states, resulting in higher performance across various tasks Bengio et al.
-(2021a); Malkin et al.
-(2022); Zhang et al.
-(2022); Jain et al.
-(2022); Deleu et al.
-(2022); Jain et al.
-(2023); Hu et al.
-(2023); Zhang et al.
-(2023); Shen et al.
-(2023b) and applications to conditioned molecule generation Shen et al.
-(2023b), maximum likelihood estimation in discrete latent variable models Hu et al.
-(2023), structure learning of Bayesian networks Deleu et al.
-(2022), scheduling computational operations Zhang et al.
-(2023), and discovering reticular materials for carbon capture Cipcigan et al.
-(2023).
-Although originally conceived for discrete state spaces, GFlowNets have been extended to more general state spaces, such as entirely continuous spaces, or spaces that are hybrid discrete-continuous Lahlou et al.
-(2023).
+Since GFlowNets learn a parametric policy, they can generalise across states, resulting in higher performance across various tasks [1, 2, 3, 4, 5, 6, 7, 8, 9] and applications to conditioned molecule generation [9], maximum likelihood estimation in discrete latent variable models [7], structure learning of Bayesian networks [5], scheduling computational operations [8], and discovering reticular materials for carbon capture [10].
+Although originally conceived for discrete state spaces, GFlowNets have been extended to more general state spaces, such as entirely continuous spaces, or spaces that are hybrid discrete-continuous [11].
 In the continuous setting, given the current state, the policy specifies a continuous probability distribution over subsequent states, and the probability density over states $x\in\mathcal{X}$ sampled with the policy is proportional to a reward density function $r(x)$.
-The continuous domain unlocks more applications for GFlowNets, such as molecular conformation sampling Volokhova et al.
-(2023) and continuous control problems Luo et al.
-(2024).
+The continuous domain unlocks more applications for GFlowNets, such as molecular conformation sampling [12] and continuous control problems [13].
 GFlowNets are trained like reinforcement learning agents.
@@
 GFlowNets therefore suffer from the same training pitfalls as reinforcement learning.
-One such issue is slow temporal credit assignment, which has thus far been addressed by designing more effective loss functions, such as detailed balance Bengio et al.
-(2021b), trajectory balance Malkin et al.
-(2022) and sub-trajectory balance Madan et al.
-(2022).
+One such issue is slow temporal credit assignment, which has thus far been addressed by designing more effective loss functions, such as detailed balance [14], trajectory balance [2] and sub-trajectory balance [15].
+This latter approach has recently been extended by providing an energy function inductive bias at the early stages of training [16].
 Besides loss functions, another aspect of GFlowNet training is the exploration strategy for acquiring training samples.
@@
 More successful strategies therefore rely on off-policy exploration.
-For the discrete setting, numerous exploration strategies have been proposed including $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al.
-(2022), Thompson sampling Rector-Brooks et al.
-(2023) and Local Search GFlowNets Kim et al.
-(2024).
-While these approaches can be generalised to the continuous domain, there is no literature benchmarking their effectiveness in this setting.
-Sampling in the continuous setting is a common occurrence in various domains such as molecular modelling Hawkins (2017); Yang et al.
-(2019) and Bayesian inference Shahriari et al.
-(2016).
+For the discrete setting, numerous exploration strategies have been proposed including $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) [17], Thompson sampling [18] and Local Search GFlowNets [19].
+While these approaches can be generalised to the continuous domain, there is limited literature benchmarking their effectiveness in this setting.
+Sampling in the continuous setting is a common occurrence in various domains such as molecular modelling [20, 21] and Bayesian inference [22].
 The local connectedness of a continuous domain allows for novel exploration strategies that are not directly applicable in the discrete setting.
-In this work, we create MetaGFN, an exploration algorithm for continuous GFlowNets inspired by metadynamics, an enhanced sampling method widely used for molecular modelling Laio and Parrinello (2002).
+[23] compared several exploration strategies in the context of diffusion samplers, and proposed alternating between on-policy sampling and backward sampling from a fixed set of MCMC samples when training continuous GFlowNets.
+These samples were selected prior to training, thereby inducing a fixed pre-training cost but keeping the cost of each training trajectory constant.
+The authors showed effective exploration, even in high-dimensional problems.
+However, their MCMC approach requires access to cheap gradients of the reward landscape, which might not always be readily available [24].
+In addition, there are some settings, such as in molecular conformation sampling, where MCMC approaches are known to require significantly longer timescales to overcome energy barriers than methods based on molecular dynamics (MD) simulations [25], thereby making the pre-training cost prohibitively expensive.
+In this work, we propose MetaGFN, a novel exploration algorithm for continuous GFlowNets inspired by metadynamics, an enhanced sampling method widely used for molecular modelling [26].
+Unlike MCMC-based sampling methods, MetaGFN operates in a general black-box setting, relying solely on an oracle for reward values, without requiring access to reward gradients.
+In common with standard metadynamics, MetaGFN converges quickly when a reduced-dimensional representation of the reward measure, specified by a so-called collective variable (CV) basis, is known.
+While our method is domain-agnostic, this feature makes MetaGFN particularly well-suited for small molecular systems, where such representations are known and often low-dimensional [27].
+MetaGFN requires no pre-training and only introduces a small, constant cost for each training trajectory, which becomes negligible as the cost of reward evaluation increases.
 The main contributions of this work are:
-  • Presenting MetaGFN, an algorithm created by adapting metadynamics to black box rewards and continuous GFlowNets;
-  • Proving that the proposed Adapted Metadynamics is consistent and reduces to standard metadynamics in a limit;
-  • Showing empirically that MetaGFN outperforms existing GFlowNets exploration strategies.
+  • Introducing MetaGFN, an algorithm that adapts metadynamics to black box rewards and continuous GFlowNets;
+  • Proving that the Adapted Metadynamics formulation underlying MetaGFN is consistent and reduces to standard metadynamics in the appropriate limit;
+  • Showing empirically that MetaGFN outperforms existing GFlowNets exploration strategies in various continuous environments, including alanine dipeptide conformation sampling.
 The rest of the paper is as follows.
-In Section 2 we review the theory of discrete and continuous GFlowNets.
-We present Adapted Metadynamics and MetaGFN in Section 3.
-In Section 4, we evaluate MetaGFN against other exploration strategies, showing that MetaGFN outperforms existing exploration strategies in two continuous environments.
+In Section 2 we review the theory of discrete and continuous GFlowNets as well as metadynamics and collective variables.
+In Section 3, we present the Adapted Metadynamics and MetaGFN algorithms.
+In Section 4, we evaluate MetaGFN against other approaches, showing that MetaGFN generally outperforms existing exploration strategies in various continuous environments.
 We finish with limitations and conclusions in Sections 5 and 6.
-Code for MetaGFN is available at [link in camera-ready].
+Code for MetaGFN is available at https://github.com/dominicp6/manifold-gfn.
 2 Preliminaries
@@
 GFlowNets also introduce additional learnable objects, such as a backward policy $P_{B}(s|s^{\prime})$, which is a distribution over the parents of any state of the DAG, to create losses that train the forward policy.
-Objective functions for GFlowNets include flow matching (FM), detailed balance (DB), trajectory balance (TB) and subtrajectory balance (STB) Bengio et al.
-(2021a, b); Malkin et al.
-(2022); Madan et al.
-(2022).
+Objective functions for GFlowNets include flow matching (FM), detailed balance (DB), trajectory balance (TB) and subtrajectory balance (STB) [1, 14, 2, 15].
 During training, the parameters of the flow objects are updated with stochastic gradients of the objective function applied to batches of trajectories.
@@
 2.2 Continuous GFlowNets
-Continuous GFlowNets extend the generative problem to continuous spaces Lahlou et al.
-(2023), where the analogous quantity to the DAG is a measurable pointed graph (MPG) Nummelin (1984).
-MPGs can model continuous spaces (e.g., Euclidean space, spheres, tori), as well as hybrid spaces, with a mix of discrete and continuous components, as often encountered in robotics, finance, and biology Neunert et al.
-(2020); Swiler et al.
-(2012); Bortolussi and Policriti (2008).
+Continuous GFlowNets extend the generative problem to continuous spaces [11], where the analogous quantity to the DAG is a measurable pointed graph (MPG) [28].
+MPGs can model continuous spaces (e.g., Euclidean space, spheres, tori), as well as hybrid spaces, with a mix of discrete and continuous components, as often encountered in robotics, finance, and biology [29, 30, 31].
 Definition 2.1 (Measurable pointed graph (MPG)).
@@
   • (iv)
-    A fully-explorable space: The number of steps required to be able to reach any measurable $B\in\Sigma$ from the source state with the forward reference kernel is bounded.
+    A fully-explorable space: The number of steps required to possibly reach any measurable $B\in\Sigma$ from the source state, and to guarantee to reach the sink state, with the forward reference kernel is bounded.
 The set of objects $(\bar{\mathcal{S}},\mathcal{T},\Sigma,s_{0},\perp,\kappa,\kappa^{b})$ then defines an MPG.
@@
 A flow F is a tuple $F=(f,P_{F})$, where $f:\Sigma\rightarrow[0,\infty)$ is a flow measure, satisfying $f(\{\perp\})=f(s_{0})=Z$, where $Z$ is the total flow.
-The reward measure is a positive and finite measure $R$ over the terminating states $\mathcal{X}$, we denote the density of this reward measure as $r(x)$, for $x\in\mathcal{X}$.
+The reward measure is a positive and finite measure $R$ over the terminating states $\mathcal{X}$.
+The density of $R$ is denoted as $r(x)$, for $x\in\mathcal{X}$.
 A flow $F$ is said to satisfy the reward-matching conditions if
 $R(dx)=f(dx)P_{F}(x,\{\perp\}).$
-If a flow satisfies the reward-matching conditions and trajectories are recursively sampled from the Markov kernel $P_{F}$ starting at $s_{0}$, the resulting measure over terminating states, $P^{\perp}(B)$, is proportional to the reward: $P^{\perp}(B)=\frac{R(B)}{R(\mathcal{X})}$ for any $B$ in the $\sigma$-algebra of terminating states Lahlou et al.
-(2023).
+If a flow satisfies the reward-matching conditions and trajectories are generated by recursively sampled from the Markov kernel $P_{F}$ starting at $s_{0}$, the resulting measure over terminating states, $P^{\perp}(B)$, is proportional to the reward: $P^{\perp}(B)=\frac{R(B)}{R(\mathcal{X})}$ for any $B$ in the $\sigma$-algebra of terminating states [11].
 Objective functions for discrete GFlowNets generalise to continuous GFlowNets.
@@
 2.3 Exploration strategies for GFlowNets
-GFlowNets can reliably learn using off-policy trajectories, a key advantage over hierarchical variational models Malkin et al.
-(2023).
-For optimal training, it is common to use a replay buffer and alternate between on-policy and off-policy (exploration) batches Shen et al.
-(2023a).
-Exploration strategies for discrete GFlowNets include $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al.
-(2022), Local Search GFlowNets Kim et al.
-(2024), and Thompson sampling (TS), which outperforms the others in grid and bit sequence domains Rector-Brooks et al.
-(2023).
-TS aims to bias exploration in regions where there is high uncertainty.
-When the forward policy is parameterised as an MLP, this is achieved using an ensemble of $K\in\mathbb{Z}^{+}$ policy heads with a common network torso.
-During training, an ensemble member is randomly sampled and used to generate a trajectory $\tau$.
-In a training batch, each ensemble member is included with probability $p$ and parameters are updated by taking a gradient step on the total loss of $\tau$ over all included members.
-No comparative literature exists on exploration strategies for continuous GFlowNets, but many methods can be adapted, and we do so here.
-For example, for TS in the continuous setting, policy heads parameterise forward policy functions instead of log probabilities.
-Another strategy unique to the continuous setting is what we call noisy exploration.
-This strategy involves introducing an additive noise parameter, denoted as $\bar{\sigma}$, to the variance parameters in the forward policy distribution, where the value of $\bar{\sigma}$ is scheduled to gradually decrease to zero over the course of training.
-2.4 Metadynamics
-Molecular dynamics (MD) uses Langevin dynamics (LD) Pavliotis (2014), a stochastic differential equation modelling particle motion with friction and random fluctuations, to simulate atomic trajectories that ergodically sample a molecule’s Gibbs measure, $\rho_{\beta}(x)\propto e^{-\beta V(x)}$.
-Here, $x$ and $p$ are atomic positions and momenta, $V(x)$ is the molecular potential, and $\beta$ is thermodynamic beta.
+GFlowNets can reliably learn using off-policy trajectories, a key advantage over hierarchical variational models [32].
+For optimal training, it is common to use a replay buffer and alternate between on-policy and off-policy (exploration) batches [33].
+In the discrete case, proposed techniques to encourage exploration include $\epsilon$-noisy exploration, tempering and the incorporation of intermediate rewards [1, 17].
+Exploration strategies for continuous GFlowNets are less well studied in the literature but several methods designed for discrete GFlowNets can be adapted [23].
+In this work, we consider:
+Local Search GFlowNets [19]: Explores by backtracking and resampling on-policy trajectories.
+Reconstructed trajectories with a higher reward than the original are used for training, thus encouraging the learning of high-reward modes.
+Thompson sampling.
+[18]: Explores high-uncertainty regions by using an ensemble of policy heads with a shared torso.
+A random head generates the on-policy trajectory, and the loss is computed by averaging contributions over heads, where each head is independently included with probability $p$.
+Noisy exploration: Explores by increasing policy uncertainty. A small constant is added to the policy variance which is gradually reduced to zero over the course of training.
+Nested sampling [34]: A Markov-Chain Monte Carlo (MCMC) algorithm is used to sample from the reward distribution.
+Backward sampling from these terminal states are used to generate off-policy trajectories.
+2.4 Metadynamics and collective Variables
+Molecular dynamics (MD) simulates that dynamics of molecules using Langevin dynamics (LD) [35], a stochastic differential equation that models particle motion under friction and random noise.
+LD trajectories ergodically sample the molecule’s Gibbs measure, $\rho_{\beta}(x)\propto e^{-\beta V(x)}$, where $x\in\mathcal{X}$ represents atomic positions, $V(x)$ is the molecular potential, and $\beta$ is the thermodynamic beta.
 [^1] We review Langevin dynamics in Appendix B.
-If the potential $V(x)$ has multiple local minima, then unbiased LD can get trapped in these minima, which can lead to inefficient sampling.
-Metadynamics is an algorithm that enhances sampling by regularly depositing repulsive Gaussian bias potentials at the centre of an evolving LD trajectory Laio and Parrinello (2002).
-At time $t>0$, metadynamics modifies LD such that the conservative component of the LD force is given by the negative gradient of the total potential, i.e.
-$-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$, where $V_{\text{bias},t}$ is the cumulative bias at time $t$.
-Bias potentials are typically chosen to only vary in the direction of user-specified, low-dimensional collective variables (CVs), $z(x):\mathcal{X}\rightarrow\mathcal{Z}$, mapping from the original space $\mathcal{X}$ to the CV space $\mathcal{Z}$.
-For biomolecules, typical CVs include protein backbone angles or distances between charge centres; these are quantities that drive and parameterise rare-event transitions.
-As the bias potential progressively fills up the potential landscape, energetic barriers are reduced, thus accelerating exploration, eventually ensuring uniform diffusion in CV space.
-In practice, the bias is defined on a regular grid, which limits CV space dimensionality to 5-10 due to exponential memory costs.
-Thus, identifying good CVs is crucial for effective metadynamics simulations in applications such as drug discovery, chemistry and materials science De Vivo et al.
-(2016); Laio and Gervasio (2008).
-In this work, we adapt the original metadynamics algorithm discussed above.
-The method has also seen numerous extensions.
-For a detailed review, see Bussi and Laio (2020) and the references therein Bussi and Laio (2020).
+However, when $V(x)$ contains multiple, deep local minima, as is common in biomolecules, then LD becomes inefficient, as it tends to get trapped in these minima, slowing exploration of the full state space.
+Metadynamics overcomes this by progressively modifying the potential landscape to discourage visits to already-explored regions [26].
+It achieves this by regularly depositing repulsive Gaussian biases, centered at the current state of the evolving LD trajectory.
+This modifies the potential to $V_{\text{total}}(x,t)=V(x)+V_{\text{bias}}(x,t)$, where $V_{\text{bias}}(x,t)$ is the cumulative bias at time $t$.
+Intuitively, metadynamics thus progressively transforms the landscape into more level surface, in which the system can diffuse more freely, thereby accelerating exploration (Figure 1).
+In the limit $t\rightarrow\infty$, the dynamics approximates free diffusion and uniformly samples the domain of $V(x)$.
+Since biases are typically specified on a numerical grid, this gives rise to a memory cost that grows exponentially with dimension.
+Therefore, biases are typically applied along low-dimensional coordinates known as collective variables (CVs).
+A collective variable $z(x):\mathcal{X}\rightarrow\mathcal{Z}$ is any mapping from the original (high-dimension) state space $\mathcal{X}$ to a lower-dimensional space $\mathcal{Z}$.
+For potential $V(x)$ and collective variables $z$, metadynamics is guaranteed to eventually uniformly sample the domain of the marginal potential $V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$.
+In molecular contexts, ideal choices of CVs correspond to the physical reaction coordinates that underly rare-event transitions (e.g.
+backbone dihedrals, interatomic distances) [36, 37].
+In more general settings, ideal CVs should resolve rare events dynamics on $V(x)$, simplfy the landscape (retaining its essential features e.g.
+barriers and basins), and minimise the dimensionality of this representation.
+For simple systems, CVs can be derived from known symmetries or macroscopic order parameters that describe state changes.
+However, if the choice of CVs is not obvious, then they can learnt through data-driven methods such as Time-Lagged Independent Component Analysis (TICA) [38], manifold learning algorithms, neural network autoencoders or variational methods [39, 40, 41].
+For an in-depth review on these CV-learning approaches, see [42].
+Metadynamics has seen numerous extensions [43].
+In the next section, we adapt the original metadynamics algorithm to the continuous black box setting.
+We will assume suitable CVs are given, but the theory we present is agnostic to their form–be it analytical expressions, neural networks, or non-parametric tabular mappings.
+Figure 1: Illustration of metadynamics in a multi-well potential. Regular deposition of a bias leads to a total potential that gradually flattens, encouraging exploration.
 3 MetaGFN: Adapted Metadynamics for GFlowNets
-Training GFlowNets in high-dimensional continuous spaces requires exploration, especially if valuable reward peaks are separated by large regions of low reward.
-In some tasks, prior knowledge of the principal manifold directions along which the reward measure varies can reduce the effective dimension required for exploration.
-This is where exploration algorithms that guarantee uniform sampling in that low-dimensional manifold, such as metadynamics, become most effective.
-With this intuition in mind, we adapt metadynamics to the black-box reward setting of continuous GFlowNets.
+Metadynamics has two properties that make it well-suited as an exploration strategy.
+Firstly, it eventually uniformly samples the domain, ensuring exploration of all local minima.
+In contrast, techniques such as Local Search GFlowNets, Thompson sampling, and noisy exploration rely on localised strategies that are prone to mode locking (Section 4).
+Secondly, its gradual diffusion from the starting configuration allows incremental exploration, resulting in more stable training than global approaches like nested sampling.
+However, metadynamics requires a potential and its gradients, which are not always available in a black-box GFlowNet setting.
+We adapt metadynamics to this setting by interpreting the reward landscape as a potential energy surface, resulting in the exploration algorithm Adapted Metadynamics (AM) (Section 3.1) and training algorithm MetaGFN (Section 3.2).
 […]
-Kernel density potential Metadynamics force computations require the gradient of the total potential, where $-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$.
-Using the above assumptions, we have $\nabla V=-\nabla r(x)/(\beta^{\prime}r(x))$.
-However, $r(x)$ is often a computationally expensive black-box function, and its gradient, $\nabla r(x)$, is unknown.
-While finite differences can estimate $\nabla r(x)$ for smooth, low-dimensional reward distributions, this approach is impractical in high-dimensional spaces.
-We avoid finite difference calculations by computing a kernel density estimate (KDE) of $V$.
-In particular, we assume collective variables $z(x)=(z_{1}(x),\dots,z_{k}(x))$, and compute both the KDE potential and bias potentials in the CV space.
-Here, each $z_{i}$ is a one-dimensional coordinate, and $z\vcentcolon\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{Z}$ is $k$-dimensional.
-As the potentials are stored on a low-dimensional grid, gradient computations are guaranteed to be cheap relative to the cost of evaluating $r(x)$.
-We use $\hat{V}(z)$ to denote the KDE estimate in CV space.
-To compute $\hat{V}(z)$, we maintain two separate KDEs: $\hat{N}(z)$ for the histogram of metadynamics states and $\hat{R}(z)$ for cumulative rewards.
-We update these KDE estimates on the fly at the same time the bias potential is updated.
-If $\mathcal{Z}\cong\mathbb{R}^{k}$, we use Gaussian kernels with update rules: [^2] If $\mathcal{Z}\cong\mathbb{T}^{k}$ ($k$-torus), we use von Mises distributions instead of Gaussians.
-$\hat{N}(z)_{t}=\hat{N}(z)_{t-1}+\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
-$\hat{R}(z)_{t}=\hat{R}(z)_{t-1}+r(x_{t})\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
-where $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\mathbb{R}^{k}$ is the kernel width, $x_{t}$ is the latest metadynamics sample and $z_{t,i}$ is the corresponding $i^{th}$ CV coordinate of $x_{t}$.
-The KDE potential $\hat{V}(x)$ is then computed as:
-$\hat{V}(x)=-\frac{1}{\beta^{\prime}}\log\left(\frac{\hat{R}(x)}{\hat{N}(x)+\epsilon}+\epsilon\right),$
-where $\epsilon>0$.
-We found empirically that $\epsilon$ ensured numerical stability by preventing division by zero and bounding the potential above by $\log\left(1/\epsilon\right)/\beta^{\prime}$.
-Taking the ratio of a reward and frequency KDE means that $\hat{V}$ rapidly and smoothly adjusts when new modes are discovered, in particular, we prove that $\hat{V}$ eventually discovers all reward modes in the CV space.
-More precisely,
-Theorem 3.1.
-If the collective variable $z(x)$ is analytic with a bounded domain, then
-$\lim_{\epsilon\rightarrow 0}\left(\lim_{\sigma\rightarrow 0}\left(\lim_{t\rightarrow\infty}\hat{V}(z,t)\right)\right)=V,$ (1)
-where $V=V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$ is the marginal potential in the CV space if $z(x)$ is not invertible.
-If $z(x)$ is invertible, $V$ is the original potential $V(x)$ in the original coordinates.
-The proof is in Appendix C.
-Implementation details We set the potential energy beta ($\beta^{\prime}$) equal to the Langevin dynamics beta ($\beta$).
-This allows a physical interpretation of $\beta$: it is inversely proportional to the (unbiased) transition rates between minima of the potential.
-[^3] From the Kramer formula; $\text{transition rate}\propto\frac{1}{\beta}\exp(\beta\Delta V)$, where $\Delta V\propto\frac{1}{\beta^{\prime}}=\frac{1}{\beta}$.
-We also set the kernel width equal to the width of the Gaussian bias.
-This is reasonable since it is the variability of $V(x)$ that determines sensible values for both these parameters.
-The resulting exploration algorithm we call Adapted Metadynamics (AM), and is presented in Algorithm 1.
-Note that the algorithm can be extended to a batch of trajectories, where each metadynamics trajectory evolves independently, but with a shared $\hat{V}$ and $V_{\text{bias}}$ which receive updates from every trajectory in the batch.
-This is the version we use in our experiments - it accelerates exploration and reduces stochastic gradient noise during training.
+3.1 Adapted Metadynamics
 Input : Manifold environment of terminating states $\mathcal{X}$ with reward density $r\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$.
 Initial state $(x_{t},p_{t})\in\mathcal{X}\times T_{x_{t}}(\mathcal{X})$.
-Collective variables $z=\left(z_{1},\dots,z_{k}\right)$.
-Parameters : Gaussian width $\sigma=\left(\sigma_{1},\dots,\sigma_{k}\right)\in\mathbb{R}^{k}$. Gaussian height $w>0$. Stride $n\in\mathbb{Z}^{+}$. LD parameters: $\gamma,\beta$. Timestep $\Delta t$.
+Collective variables $z=\left(z_{1},\dots,z_{d}\right)$.
+Parameters : Gaussian width $\sigma=\left(\sigma_{1},\dots,\sigma_{d}\right)\in\mathbb{R}^{d}$. Gaussian height $w>0$. Stride $n\in\mathbb{Z}^{+}$. LD parameters: $\gamma,\beta$. Timestep $\Delta t$.
 1
@@
 8   every n timesteps $n\Delta t$:
-9     $\hat{N}\leftarrow\hat{N}+\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
-10     $\hat{R}\leftarrow\hat{R}+r(x_{t})\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
+9     $\hat{N}\leftarrow\hat{N}+\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
+10     $\hat{R}\leftarrow\hat{R}+r(x_{t})\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
 11     $\hat{V}\leftarrow-\frac{1}{\beta}\log\left(\hat{R}/(\hat{N}+\epsilon)+\epsilon\right)$
-12     $V_{\text{bias}}(z)\leftarrow V_{\text{bias}}(z)+n\cdot\Delta t\cdot w\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-{z}_{t,i}}{{\sigma_{i}}}\right|^{2}}\right)$
+12     $V_{\text{bias}}(z)\leftarrow V_{\text{bias}}(z)+n\cdot\Delta t\cdot w\cdot\exp\left({-\frac{1}{2}\sum_{i=1}^{d}\frac{(z_{i}-{z}_{i,t})^{2}}{{\sigma_{i}^{2}}}}\right)$
 13   compute forces:
 14     $F\leftarrow-\left(\nabla_{z}\hat{V}(z)|_{z=z_{t}}+\nabla_{z}V_{\text{bias}}(z)|_{z=z_{t}}\right)\cdot\nabla_{x}z|_{x=x_{t}}$
-  propagate $x_{t},p_{t}$ by $\Delta t$ using Langevin dynamics with computed force $F$ (Alg. 2, Appendix D).
+  propagate $x_{t},p_{t}$ by $\Delta t$ using Langevin dynamics with computed force $F$ (Alg. 3, Appendix B).
 Algorithm 1 Adapted Metadynamics
-Training GFlowNets with Adaptive Metadynamics (MetaGFN) Each Adaptive Metadynamics sample $x_{i}\in\mathcal{X}$ is an off-policy terminal state sample.
+Metadynamics requires the gradient of the total potential, where $-\nabla V_{\text{total}}(x,t)=-\nabla(V(x)+V_{\text{bias}}(x,t))$.
+Using the above assumptions, we have $\nabla V(x)=-\nabla r(x)/(\beta^{\prime}r(x))$.
+However, $r(x)$ is often a computationally expensive black-box function, and its gradient, $\nabla r(x)$, is unknown.
+While finite differences can estimate $\nabla r(x)$ for smooth, low-dimensional reward distributions, this approach is impractical in high-dimensional spaces.
+Below, we explain how to adapt metadynamics to this black-box setting and avoid finite difference gradient estimates by storing a dynamically-updated kernel density estimate (KDE) of the potential.
+We call the modified metadynamics algorithm Adapted Metadynamics (Algorithm 1).
+Let $x_{t}$ denote the metadynamics sample at time $t$ and $z(x)=(z_{1}(x),\dots,z_{d}(x))$ be a given set of collective variables, where $z_{i}$ is a one-dimensional coordinate, and $z\in\mathcal{Z}$ is $d$-dimensional.
+Let $z_{i,t}\vcentcolon=z_{i}(x_{t})$ denote the corresponding $i^{th}$ CV coordinate at time $t$ and assume that the Jacobian $\nabla_{x}z$ is well-defined.
+We store a discretisation of the KDE marginal potential and bias potentials in CV space $\mathcal{Z}$.
+Since these are stored in memory, gradient computations are cheap and require no further evaluations of $r(x)$.
+Let $\hat{V}(z,t)$ represent the KDE of the marginal potential at time $t$.
+To compute $\hat{V}(z,t)$, we maintain two separate KDEs: $\hat{N}(z,t)$ for visited states (line 1) and $\hat{R}(z,t)$ for cumulative rewards (line 1).
+We update these KDEs on-the-fly at the same time the bias potential is updated, which occurs every integer $n$ steps of Langevin dynamics (line 1).
+If $\mathcal{Z}\cong\mathbb{R}^{d}$, we use Gaussian kernels with kernel width $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\mathbb{R}^{k}$, matching the width of the Gaussian bias (line 1) [^3] If $\mathcal{Z}\cong\mathbb{T}^{d}$ ($d$-torus), we use von Mises distributions..
+This is a reasonable as both are set by the variability length scale of $V(x)$.
+Finally, the KDE potential $\hat{V}(z,t)$ is then computed as $\hat{V}(z,t)=-\frac{1}{\beta}\log\left(\frac{\hat{R}(z,t)}{\hat{N}(z,t)+\epsilon}+\epsilon\right)$, for a fixed constant $\epsilon>0$ (line 1).
+We found empirically that $\epsilon$ ensured numerical stability by preventing division by zero and bounding $\hat{V}$ from above.
+Defining $\hat{V}$ through the ratio $\hat{R}/\hat{N}$ ensures that it rapidly and smoothly adjusts whenever new modes are discovered.
+Furthermore, we prove that $\hat{V}$ eventually discovers all reward modes in the CV space.
+More precisely,
+Theorem 3.1.
+If the collective variable $z(x)$ is analytic with a bounded domain, then
+$\lim_{\epsilon\rightarrow 0}\left(\lim_{\sigma\rightarrow 0}\left(\lim_{t\rightarrow\infty}\hat{V}(z,t)\right)\right)=V,$ (1)
+where $V=V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$.
+The proof is in Appendix C.
+Note that the algorithm can be extended to a batch of trajectories, where each metadynamics trajectory evolves independently, but with a shared $\hat{V}$ and $V_{\text{bias}}$ which receive updates from every trajectory in the batch.
+We found empirically that this accelerates exploration and reduces stochastic gradient noise during training and this is the version we use in our experiments (Section 4).
+3.2 MetaGFN
+Each Adapted Metadynamics sample $x_{i}\in\mathcal{X}$ is an off-policy terminal state sample.
 To train a GFlowNet, complete trajectories are required.
@@
 However, since this requires a backward policy, this is compatible with DB, TB, and STB losses, but not FM loss.
-Given the superior credit assignment of the former losses, this is not a limitation Madan et al.
-(2022).
+Given the superior credit assignment of the former losses, this is not a limitation [15].
 Additionally, we use a replay buffer.
-Due to the theoretical guarantee that Adaptive Metadynamics will eventually sample all collective variable space (Theorem 3.1), AM samples are ideal candidates for storing in a replay buffer.
+Due to the theoretical guarantee that Adapted Metadynamics will eventually sample all collective variable space (Theorem 3.1), AM samples are ideal candidates for storing in a replay buffer.
 When storing these trajectories in the replay buffer, there are two obvious choices:
@@
   2. 2.
-    Store only the Adaptive Metadynamics sample and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
-We investigate both options in our experiments. We call the overall training algorithm MetaGFN, with pseudocode presented in Algorithm 3, Appendix D.
+    Store only the Adapted Metadynamics sample and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
+We investigated both options in preliminary experiments (Appendix D.2).
+Option 2 was found to be uniformly superior and it is the version we use in our main experiments (Section 4).
+We call the overall training algorithm MetaGFN, with pseudocode presented in Algorithm 2, below.
+Input : Forward policy $P_{F}$. Backwards policy $P_{B}$. Loss function $L$.
+Parameters : How often to run Adapted Metadynamics batches, freqMD. How often to run replay buffer batches, freqRB. Batch size, $b$. Stride, $n\in\mathbb{Z}^{+}$. Time step, $\Delta t>0$
+1
+2for each episode do:
+3if episode number is divisible by freqMD:
+4   Run Adapted Metadynamics (batch size $b$) for time $n\Delta t$, obtain samples $\{x_{1},\dots,x_{b}\}$
+5   Push $\{x_{1},\dots,x_{b}\}$ to the replay buffer
+6   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
+7elif episode number is divisible by freqRB:
+8   Random sample $\{x_{1},\dots,x_{b}\}$ from the replay buffer
+9   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
+10
+11else:
+12   Generate trajectories $\{\tau_{1},\dots,\tau_{b}\}$ on-policy
+13
+14 Compute loss $l=\sum_{i=1}^{b}L(\tau_{i},P_{F},P_{B})$
+ Take gradient step on loss $l$
+Algorithm 2 MetaGFN
 4 Experiments
-4.1 Line environment
-Figure 1: Line Environment reward function, equation (2).
-We consider a one-dimensional line environment, with state space $\mathcal{S}=\mathbb{R}\times\left\{t\in\mathbb{N},1\leq t\leq 3\right\}$, where $t$ indexes the position of a state in a trajectory.
-The source state is $s_{0}=(0,0)$ and trajectories terminate after exactly $3$ steps.
+We compare MetaGFN with Thompson sampling, noisy, Local Search GFlowNets and nested sampling (Section 2.3).
+We run experiments in five continuous environments, summarised below.
+For each exploration strategy, we use a replay buffer and alternate between exploration and replay buffer batches.
+For MetaGFN, we use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$.
+Forward and backward kernels are Gaussian/von Mises mixture distributions, with distribution parameters specified at each state by an MLP.
+In all environments, the additional computational expense of running Adapted Metadynamics was negligible (<5%) compared to the training time of the models.
+Full experimental details are in Appendix D.
+Figure 2: Line environment reward density.
+Line environment: A one-dimensional environment with state space $\mathcal{S}=\mathbb{R}\times\left\{t\in\mathbb{N},1\leq t\leq 3\right\}$, where $t$ indexes the position of a state in a trajectory.
+The source state is $s_{0}=(0,0)$.
 The terminal states are therefore $\mathcal{X}=\mathbb{R}\times\left\{3\right\}\cong\mathbb{R}$.
+The collective variable is the identity, i.e.
+$z=x$.
 The reward density, plotted in Figure 2, consists of an asymmetric bimodal peak near the origin and an additional distant lone peak.
@@
 where $\mathcal{N}(\mu,\sigma^{2})$ is a Gaussian density with mean $\mu$ and variance $\sigma^{2}$.
-The forward and backward probability transition kernels are a mixture of $3$ Gaussian distributions which, along with the flow $\hat{f}$, are parameterised by an MLP.
-We compare the following exploration strategies: entirely on-policy (no exploration), noisy exploration, Thompson Sampling, and MetaGFN.
-For each strategy, we use a replay buffer and alternate between an exploration batch and a replay buffer batch.
-For MetaGFN, we use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$ and always regenerate MetaGFN training trajectories using the current backward policy.
-As the environment is one-dimensional, the collective variable is simply $z(x)=x$.
-We evaluate performance by computing the L1 error between the known reward distribution and the empirical on-policy distribution during training (see Appendix E.1 for the full experimental details).
-Figure 2: The L1 difference between on-policy and reward distribution during training in the line environment for different loss functions and exploration strategies.
+Figure 3: Free energy surface of alanine dipeptide.
+Alanine dipeptide environment: One application of continuous GFlowNets is molecular conformation sampling [12].
+Here, we train a GFlowNet to sample conformational states of alanine dipeptide (AD), a a small biomolecule of 23 atoms that plays a key role in modelling the dynamics of proteins [44].
+The metastable states of AD are distinguished in a 2D CV space defined by the backbone dihedral angles $\phi$ and $\psi$.
+The resulting free energy surface $V(\phi,\psi)$ in explicit water, obtained after extensive sampling with long molecular dynamics simulations, is shown in Figure 3.
+The metastable states, in increasing energy, are $P_{||}$, $\alpha_{R}$, $C_{5}$, $\alpha^{\prime}$, $\alpha_{L}$, and $\alpha_{D}$.
+The state space is $\mathcal{S}=\mathbb{T}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$, with source state $s_{0}=P_{||}=(-1.2,2.68)$.
+Terminal states are $\mathcal{X}=\mathbb{T}^{2}\times\{3\}\cong\mathbb{T}^{2}$.
+The reward function is the Boltzmann weight, $r(\phi,\psi)=\frac{1}{Z}\exp(-\beta V(\phi,\psi))$, where $Z$ is the normalisation constant.
+Figure 4: Grid environment reward density in dimension 2.
+Grid environments: We consider (hyper)grids in $d=2$, $3$ and $4$ dimensions.
+The state space is $\mathcal{S}=[-15,15]^{d}\times\{t\in\mathbb{N},1\leq t\leq 3\}$.
+The $k$-dimensional (hyper)grid consists of $2^{k}$ modes, located at the corners of the (hyper)cube $[-10,10]^{d}$.
+The (hyper)grid is centered on the origin with an edge width of $20$.
+The reward modes are Gaussians with variance $\sigma^{2}=2$ (Figure 4).
+We use trajectory lengths of 3, 5, and 6 for dimensions 2, 3, and 4 respectively.
+The collective variable is the identity, i.e.
+$z=x$.
+The source state is the origin and terminal states are in $\mathcal{X}\cong\mathbb{R}^{k}$.
+4.1 Results
+In each environment, we run experiments with three different loss functions: Detailed Balance (DB), Trajectory Balance (TB) and Subtrajectory Balance (STB).
+We evaluate performance by computing the L1 error between the known reward distribution and the empirical on-policy distribution during training resulting from $10^{4}$ independent samples.
+The results, averaged over $10$ repeats, are shown in Figure 5.
+In 12 out of 15 of the experiments, MetaGFN outperformed all other exploration strategies.
+We give further analysis below.
+Line Environment: Among the losses, TB shows the lowest variance, and for all losses, MetaGFN always converges to a lowest error  [^4] We do not show results for nested sampling on the line environment as the implementation used did not support one-dimensional environments, see Appendix D..
+Indeed, MetaGFN is the only method that consistently samples the distant reward peak at $x=20$ (Appendix D.2).
+Although other strategies occasionally sampled the distant peak, MetaGFN alone converges because it continuously samples this peak, even if the forward policy starts to lock onto the central modes, thus ensuring that the replay buffer is always populated with diverse samples.
+The slight increase in the loss of MetaGFN around batch number $5\times 10^{3}$ occurs as the on-policy distribution widens when Adapted Metadynamics first discovers the distant peak.
+Appendix D.2 provides further analysis of Adapted Metadynamics and compares different MetaGFN variants, with and without noise, and with and without trajectory regeneration.
+We confirm that the variant of MetaGFN presented in Figure 5 (no added noise, always regenerate trajectories) is the most robust.
+Alanine Dipeptide Environment: For each loss, MetaGFN generally converges to a lower error than all other exploration strategies.
+However, for TB loss, the average L1 error is marginally higher than on-policy training, but this conceals the fact that the best-case error is smaller.
+To better understand this result, we examined the best and worst training runs (as measured by L1 error) for TB on-policy and TB MetaGFN, shown in Figure 6.
+Note that, unlike on-policy, the best MetaGFN run learns to sample the rare $\alpha_{L}$ mode.
+In the worst case, MetaGFN fails to converge (although this is rare; only one of 10 runs failed).
+In Table 1, we quantify how often the different AD modes were sampled over the different repeats (a mode is considered sampled if the on-policy distribution has a mode within the correct basin of attraction).
+TB loss with MetaGFN is the only combination that consistently samples the majority of modes.
+The only mode not sampled by any method is $\alpha_{D}$, which has a natural abundance approximately $10$ times less frequent than $\alpha_{L}$.
+Figure 5: The L1 difference between on-policy and reward distribution during training for different loss functions and exploration strategies.
 The mean is plotted with standard error over 10 repeats.
 DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
-The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 2.
-Among the three loss functions, TB loss has the lowest variance loss profiles, and MetaGFN consistently converges to a lower minimum error than all other exploration strategies.
-Indeed, MetaGFN was the only method that consistently sampled the distant reward peak at $x=20$, while other methods plateau in error after locking onto the central modes (Appendix E.2).
-Despite this, we observed that in the initial stages of training, all exploration strategies found occasional samples from the distant peak.
-The reason MetaGFN is the only method that converges is that Adaptive Metadynamics manages to consistently sample the distant peak, even if the on-policy starts to focus on the central modes.
-This keeps the replay buffer populated with samples from every reward peak during training, which eventually encourages the on-policy to sample from every mode.
-The small increase in the loss of MetaGFN around batch number $5\times 10^{3}$ happens because the on-policy distribution widens when Adaptive Metadynamics first discovers the distant peak.
-In Appendix E.2, we show further details of Adaptive Metadynamics in this environment and we compare different MetaGFN variants, with and without noise, and with and without trajectory regeneration.
-We confirm that the version of MetaGFN presented in Figure 2 (no added noise and always regenerate trajectories) is the most robust variant.
-4.2 Grid environment
-Figure 3: Grid Environment reward function.
-We consider a two-dimensional environment, with state space $\mathcal{S}=\mathbb{R}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$.
-The source state is $s_{0}=(0,0,0)$ and trajectories terminate after exactly $3$ steps.
-Terminal states are $\mathcal{X}=\mathbb{R}^{2}\times\{3\}\cong\mathbb{R}^{2}$.
-The reward function is given by the sum of four symmetric Gaussians with means $\mu=(-7,-7),(-7,7),(7,7),(7,-7)$ and variances $\sigma^{2}=2$ (Figure 3).
-The forward and backward probability transition kernels are a mixture of $4$ bivariate Gaussian distributions with means and diagonal variance matrices, along with the flow $\hat{f}$, parameterised by an MLP.
-We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix F for full experiment details).
-Figure 4: The L1 difference between on-policy and reward distribution during training in the grid environment for different loss functions and exploration strategies.
-The mean is plotted with standard error over 10 repeats.
-DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
-The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 4.
-We see that for most loss functions and exploration strategies (except Thompson sampling with TB loss), the training curves converge to a low loss, indicating that all four modes of the grid environment are well-sampled.
-However, in all cases, MetaGFN converges faster than other exploration strategies.
-The fastest convergence occurs for MetaGFN with TB loss; after $10^{3}$ batches MetaGFN has reached a similar loss that noisy exploration reaches after $10^{4}$ batches and on-policy reaches after $\sim 3\times 10^{4}$ batches.
-Meanwhile, the loss curves have lower error indicating more stable training with MetaGFN in this environment.
-4.3 Alanine dipeptide environment
-Figure 5: (a) Free energy surface of alanine dipeptide in an explicit solvent with all metastable states annotated.
-(b) The potential KDE $\hat{V}$ learnt using Adapted Metadynamics with von Mises kernel, concentration $\kappa=10$.
-One application of continuous GFlowNets is molecular conformation sampling Volokhova et al.
-(2023).
-Here, we train a GFlowNet to sample conformational states of alanine dipeptide (AD), a small biomolecule of $23$ atoms that plays a key role in modelling backbone dynamics of proteins Hermans (2011).
-The metastable states of AD can be distinguished in a two-dimensional CV space of $\phi$ and $\psi$, the two backbone dihedral angles.
-The resulting free energy surface for AD in explicit water, $V(\phi,\psi)$, obtained after extensive sampling long molecular dynamics simulation is shown in Figure 5(a).
-The metastable states (energy minima), in increasing energy, are $P_{||}$, $\alpha_{R}$, $C_{5}$, $\alpha^{\prime}$, $\alpha_{L}$ and $\alpha_{D}$.
-The state space is $\mathcal{S}=\mathbb{T}^{2}\times\{t\in\mathbb{N},1\leq t\leq 3\}$.
-The source state is $s_{0}=P_{||}=(-1.2,2.68)$ and trajectories terminate after exactly $3$ steps.
-Terminal states are $\mathcal{X}=\mathbb{T}^{2}\times\{3\}\cong\mathbb{T}^{2}$.
-The reward density is given by the Boltzmann weight, $r(\phi,\psi)=\frac{1}{Z}\exp(-\beta V(\phi,\psi))$, where $Z$ is the normalisation constant.
-The forward and backward probability transition kernels are defined as a mixture of $3$ bivariate von Mises distributions, parameterised through an MLP.
-We consider the same exploration strategies and evaluation measures as for the Line Environment (see Appendix G for full experimental details).
-Figure 6: The L1 difference between on-policy and reward distribution during training in the alanine dipeptide environment for different loss functions and exploration strategies.
-The mean is plotted with a standard error over 10 repeats.
-DB - Detailed Balance loss, TB - Trajectory Balance loss, STB - Subtrajectory Balance loss.
-Figure 7: Learned on-policy distribution for TB on-policy and TB MetaGFN training runs.
-The colour bar shows the probability density.
-Red histograms show the marginal distribution along the angular coordinates.
-Black curves show the marginal distributions of the ground truth.
-In the best case, MetaGFN is able to learn the $\alpha_{L}$ mode.
-In the worst case, MetaGFN fails to converge.
-On-policy training, although more consistent, fails to learn to sample from the $\alpha_{L}$ mode.
-The results (mean and standard deviation over $10$ random seeds) for each loss function and exploration strategy are shown in Figure 6.
-For each loss function, models trained with MetaGFN generally converge to a lower minimum error than all other exploration strategies.
-For TB loss, however, the average L1 error is marginally higher than on-policy training, but this conceals the fact that the best-case error is smaller.
-Thus, to better understand this result, we examine the best and worst training runs (as measured by L1 error) for TB on-policy and TB MetaGFN, shown in Figure 7.
-We see that the best run trained with metadynamics can sample from the rare $\alpha_{L}$ minima, unlike the on-policy run.
-In the worst run, MetaGFN fails to converge (although this is rare; only one of 10 runs failed).
-In Table 1, we quantify how often the different AD modes were correctly sampled over the different repeats (a mode is correctly sampled if the on-policy distribution also has a mode within the correct basin of attraction).
-The only mode not correctly sampled by any method is $\alpha_{D}$, which has a natural abundance approximately $10$ times less frequent than $\alpha_{L}$.
-We see that TB loss with MetaGFN is the only combination that can consistently sample the majority of modes, whilst noisy exploration and Thompson Sampling both perform worse than on-policy in this environment.
 Table 1: Number of correct samples of AD modes in trained GFlowNets over 10 independent repeats for DB, STB, and TB loss functions.
@@
 $\alpha_{L}$ 0 1 1 0 0 8
+Figure 6: Learned on-policy distribution for TB on-policy and TB MetaGFN training runs.
+The colour bar shows the probability density.
+Red histograms show the marginal distribution along the angular coordinates.
+Black curves show the marginal distributions of the ground truth.
+In the best case, MetaGFN is able to learn the $\alpha_{L}$ mode.
+In the worst case, MetaGFN fails to converge.
+On-policy training, although more consistent, fails to learn to sample from the $\alpha_{L}$ mode.
+Grid Environments: MetaGFN achieves the lowest error in 7 of 9 experiments, although note that the performance of all methods generally decreases with increasing dimension.
+This is unsurprising, since the learning task becomes more difficult.
+For MetaGFN, this is also an expected outcome of the curse of dimensionality of the replay buffer grid; in high-dimensional spaces metadynamics samples encounter high-reward modes less frequently, leading to a replay buffer with less sample diversity.
+This is a known limitation of metadynamics [26].
+Nested Sampling outperforms MetaGFN in one case (Grid 3D, DB loss) but relies on costly MCMC-based pre-training that scales exponentially with dimension and requires reward gradients, unlike MetaGFN.
+However, the generally favourable performance of MetaGFN could be due to the method discovering an overall larger diversity of off-policy samples, that are refreshed and updated during training, in contrast to the static samples when using Nested Sampling approach.
 5 Limitations
-For metadynamics to be an effective sampler, the CVs must be low-dimensional and bounded, properties that were satisfied in both our experimental environments.
-Therefore, it is necessary to either know such CVs in advance, assuming they exist or learn them automatically from data Sidky et al.
-(2020).
-An alternative approach would be to learn CVs adaptively by parameterising the CV function by a neural network and updating its parameters by back-propagating through the GFlowNet loss when training on MetaGFN trajectories.
-A final improvement could be to replace the metadynamics algorithm itself with a variant with smoother convergence properties, such as well-tempered metadynamics Barducci et al.
-(2008) or on-the-fly probability enhanced sampling (OPES) Invernizzi (2021).
-We leave these extensions for future work.
+For metadynamics to be an effective sampler, the CVs must be low-dimensional and bounded.
+Either these CVs should be known analytically (as in our experiments) or learnt from data [42].
+Alternatively, CVs could be learnt adaptively by parameterising $z(x;\theta)$ by a neural network and updating its parameters by back-propagating through the GFlowNet loss during training.
+The metadynamics algorithm could also be replaced with a variant with smoother convergence properties, such as well-tempered metadynamics [45] or on-the-fly probability enhanced sampling (OPES) [46].
+We leave these as extensions for future work.
 6 Conclusions
@@
 Importantly, this work advocates an approach wherein techniques derived from molecular modelling can be adapted for machine learning tasks.
-Looking ahead, we anticipate that this could be a fruitful area of cross-disciplinary research, where existing ideas from the enhanced sampling literature can find applications in a broad range of generative modelling and reinforcement learning tasks.
+Looking ahead, we anticipate that this could be a fruitful area of cross-disciplinary research, where existing ideas from the enhanced sampling literature can find further applications in generative modelling and reinforcement learning.
 Appendix A Loss functions
@@
 In the above, $x,p\in\mathbb{R}^{D}$ are vectors of instantaneous position and momenta respectively, $F\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ is a force function, $W(t)$ is a vector of $D$ independent Wiener processes, $M\in\mathbb{R}^{D}\times\mathbb{R}^{D}$ is a constant diagonal mass matrix, and $\gamma,\beta>0$ are constant scalars which can be interpreted as a friction coefficient and inverse temperature respectively.
-In conventional Langevin dynamics, the force function is given by the gradient of the potential energy function, $F=-\nabla V$, where $V\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the dynamics are ergodic with respect to the Gibbs-Boltzmann density
+In conventional Langevin dynamics, the force function is given by the gradient of the potential energy function, $F(x)=-\nabla V(x)$, where $V\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the dynamics are ergodic with respect to the Gibbs-Boltzmann density
 $\rho_{\beta}(x,p)\propto e^{-\beta H(x,p)},$
 where $H(x,p)=p^{T}M^{-1}p/2+V(x)$ is the Hamiltonian.
-Since the Hamiltonian is separable in position and momenta terms, the marginal Gibbs-Boltzmann density is position space is simply $\rho_{\beta}(x)\propto e^{-\beta V(x)}$.
+Since the Hamiltonian is separable in position and momenta terms, the marginal Gibbs-Boltzmann density is position space is simply $\rho_{\beta}(x)\propto e^{-\beta V(x)}$, often refered to as the Gibbs density.
+We present the Euler-Maruyama numerican scheme for integrating equations 3 below. This is called in line 1 of Adapted Metadynamics (Algorithm 1).
+Input : Current state $(x_{t},p_{t})$. Force $F$.
+Parameters : Friction coefficient $\gamma$. Thermodynamic beta $\beta$. Timestep $\Delta t$.
+Output : State $(x_{t+\Delta t},p_{t+\Delta t})$ at the next timestep.
+1
+2Sample a random vector $R$, with the same dimension as $x_{t}$, where each element is an independent sample from a standard normal.
+3 $x_{t+\Delta t}=x_{t}+p_{t}\Delta t$
+4 $p_{t+\Delta t}=p_{t}+F\Delta t-\gamma p_{t}\Delta t+\sqrt{2\gamma\Delta t/\beta}\cdot R$
+5return $(x_{t+\Delta t},p_{t+\Delta t})$
+Algorithm 3 Euler-Maruyama Langevin Dynamics Step
 Appendix C Proofs
@@
 For concreteness, throughout this proof we assume that the kernel function is a Gaussian. We explain at the appropriate stage in the proof, indicated by (*), how this assumption can be relaxed.
-First we take the $t\rightarrow\infty$ limit. Since the log function is continuous, the limit and log can be interchanged and we have
-$\lim_{t\rightarrow\infty}\hat{V}(x,t)=-\frac{1}{\beta^{\prime}}\log\left(\lim_{t\rightarrow\infty}\left(\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}\right)+\epsilon\right).$
-Recall the uniform time discretisation $t_{n}=n\Delta t$ of metadynamics (Algorithm 1). Thus, we can write
-$\hat{R}(x,t_{n})=\sum_{i=1}^{n}r(x_{i})\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)},$ (8)
-$\hat{N}(x,t_{n})=\sum_{i=1}^{n}\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$ (9)
-Since the domain is bounded, we know that for fixed $x$, both (8) and (9) have limit at infinity, i.e.
-$\lim_{t\rightarrow\infty}\hat{R}(x,t)=\infty$ and $\lim_{t\rightarrow\infty}\hat{N}(x,t)=\infty$.
+First we take the $n\rightarrow\infty$ limit.
+Recall the notation from Section 3, i.e.
+assume collective variables $z(x)=(z_{1}(x),\dots,z_{d}(x))$ where $z\vcentcolon\mathcal{X}\rightarrow\mathcal{Z}$, and $\mathcal{Z}$ is $d$-dimensional.
+Since the log function is continuous, the limit and log can be interchanged and we have
+$\lim_{n\rightarrow\infty}\hat{V}(z,t_{n})=-\frac{1}{\beta^{\prime}}\log\left(\lim_{n\rightarrow\infty}\left(\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})+\epsilon}\right)+\epsilon\right),$
+where, from Algorithm 1:
+$\hat{R}(z,t_{n})=\sum_{i=1}^{n}r(x_{t_{i}})\exp{\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)},$ (8)
+$\hat{N}(z,t_{n})=\sum_{i=1}^{n}\exp{\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$ (9)
+Since the domain is bounded, we know that for fixed $z$, both (8) and (9) have limit at infinity, i.e.
+$\lim_{n\rightarrow\infty}\hat{R}(z,t_{n})=\infty$ and $\lim_{n\rightarrow\infty}\hat{N}(z,t_{n})=\infty$.
 Hence, by Lemma C.1, we have
-$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}=\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)},$
+$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})+\epsilon}=\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})},$
 provided the limit on the RHS exists. Next, we show that this limit exists by computing it explicitly. The limit can be written
-$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)}=\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(x_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
+$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(x_{t_{i}})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$
 Recall that metadynamics eventually leads to uniform sampling over the domain, independent of the potential.
-Hence, since $R$ and $z(x)$ are analytic, by Lemma C.2 we may replace the metadynamics samples $x_{i}$ with samples from a uniform distribution, denoted $u_{i}$:
-$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(u_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
+Hence, since $R$ and $z(x)$ are analytic, by Lemma C.2 we may replace the metadynamics samples $x_{t_{i}}$ with samples $u_{i}$ from a uniform distribution over $\mathcal{X}$:
+$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(u_{i})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(u_{i}))^{2}}{2{\sigma}_{j}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(u_{i}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$
 In the limit, the ratio of sums with uniform sampling becomes a ratio of integrals:
-$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\frac{\int_{D}r(x^{\prime})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}}{\int_{D}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}},$
-where $D\subset\mathbb{R}^{d}$ is the domain.
+$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\frac{\int_{\mathcal{X}}r(x^{\prime})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x^{\prime}))^{2}}{2{\sigma}_{j}^{2}}\right)\mathrm{d}x^{\prime}}{\int_{\mathcal{X}}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x^{\prime}))^{2}}{2{\sigma}_{j}^{2}}\right)\mathrm{d}x^{\prime}}.$
 The limit is therefore a (scaled) convolution of the reward function with a Gaussian in the collective variable space with width vector $\sigma$.
-Taking the limit $\sigma_{i}\rightarrow 0$ for all $i\in\left\{1,2,\dots,d\right\}$, the Gaussian convergences to a delta distribution in the collective variable space and we have
-$\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}.$
+Taking the limit $\sigma_{j}\rightarrow 0$ for all $j\in\left\{1,2,\dots,d\right\}$, the Gaussian convergences to a delta distribution in the collective variable space and we have
+$\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\int_{\mathcal{X}}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}.$
 (*) This step also holds for any kernel that becomes distributionally equivalent to a Dirac delta function in the limit that its variance parameter goes to zero.
@@
 Finally, we take the limit $\epsilon\rightarrow 0$ to obtain
-$\lim_{\epsilon\rightarrow 0}\lim_{\sigma\rightarrow 0}\lim_{t\rightarrow\infty}\hat{V}(x,t)$ $=-\frac{1}{\beta^{\prime}}\lim_{\epsilon\rightarrow 0}\log\left(\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}+\epsilon\right)$ (10)
-$=-\frac{1}{\beta^{\prime}}\int_{D}\log\left(r(x^{\prime})\right)\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}$ (11)
-$=\int_{D}V(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}\vcentcolon=V(z),$ (12)
+$\lim_{\epsilon\rightarrow 0}\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\hat{V}(z,t_{n})$ $=-\frac{1}{\beta^{\prime}}\lim_{\epsilon\rightarrow 0}\log\left(\int_{\mathcal{X}}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}+\epsilon\right)$ (10)
+$=-\frac{1}{\beta^{\prime}}\int_{\mathcal{X}}\log\left(r(x^{\prime})\right)\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}$ (11)
+$=\int_{\mathcal{X}}V(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}\vcentcolon=V(z),$ (12)
 where we have used the definition $V(x^{\prime})=-\frac{1}{\beta^{\prime}}\log(r(x^{\prime}))$ and in the last step we used the definition of the marginal potential energy in the collective variable space.
@@
 ∎
-Appendix D Algorithms
-Input : Current state $(x_{t},p_{t})$. Force $F$.
-Parameters : Friction coefficient $\gamma$. Thermodynamic beta $\beta$. Timestep $\Delta t$.
-Output : State $(x_{t+\Delta t},p_{t+\Delta t})$ at the next timestep.
-1
-2Sample a random vector $R$, with the same dimension as $x_{t}$, where each element is an independent sample from a standard normal.
-3 $x_{t+\Delta t}=x_{t}+p_{t}\Delta t$
-4 $p_{t+\Delta t}=p_{t}+F\Delta t-\gamma p_{t}\Delta t+\sqrt{2\gamma\Delta t/\beta}\cdot R$
-5return $(x_{t+\Delta t},p_{t+\Delta t})$
-Algorithm 2 Euler-Maruyama Langevin Dynamics Step
-In the algorithm below, we present the variant of MetaGFN where we store Adaptive Metadynamics samples in the replay buffer and regenerate trajectories using the current backward policy when retrieving from the replay buffer.
-This is the variant we used in our experiments in the main text.
-Input : Forward policy $P_{F}$. Backwards policy $P_{B}$. Loss function $L$.
-Parameters : How often to run Adaptive Metadynamics batches, freqMD. How often to run replay buffer batches, freqRB. Batch size, $b$. Stride, $n\in\mathbb{Z}^{+}$. Time step, $\Delta t>0$
-1
-2for each episode do:
-3if episode number is divisible by freqMD:
-4   Run Adaptive Metadynamics (batch size $b$) for time $n\Delta t$, obtain samples $\{x_{1},\dots,x_{b}\}$
-5   Push $\{x_{1},\dots,x_{b}\}$ to the replay buffer
-6   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
-7elif episode number is divisible by freqRB:
-8   Random sample $\{x_{1},\dots,x_{b}\}$ from the replay buffer
-9   Backward sample from $\{x_{1},\dots,x_{b}\}$ using current $P_{B}$ to obtain trajectories $\{\tau_{1},\dots,\tau_{b}\}$
-10
-11else:
-12   Generate trajectories $\{\tau_{1},\dots,\tau_{b}\}$ on-policy
-13
-14 Compute loss $l=\sum_{i=1}^{b}L(\tau_{i},P_{F},P_{B})$
- Take gradient step on loss $l$
-Algorithm 3 MetaGFN
-Appendix E Experiment details: line environment
-E.1 Experimental setup
-Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 256 hidden units per layer.
+Appendix D Experiment details
+D.1 Common experimental parameters
+Training parameters: For all environments we train for $B=10^{5}$ batches use a learning rate with a linear schedule, starting at $10^{-3}$ and finishing at 00.
+For the TB loss, we train the $\log Z_{\theta}$ with a higher initial learning rate of $10^{-1}$ (also linearly scheduled).
+For the STB loss, we use $\lambda=0.9$, a value that has worked well in the discrete setting [15].
+We use the Adam optimiser with gradient clipping.
+For all loss functions, we clip the minimum log-reward signal at $-10$.
+This enables the model to learn despite regions of near-zero reward between the modes of $r(x)$.
+Replay buffer: We use a replay buffer with capacity for $10^{4}$ trajectories.
+Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-3}$.
+When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
+Figure 7: Exponential noise schedule.
+Noisy exploration: An additional constant, $\bar{\sigma}$, is added to the standard deviations of the Gaussian distributions of the forward and backward policies.
+Specifically, the forward policy becomes $\hat{p}_{F}(s_{t},s_{t-1};\theta)=\sum_{i=1}^{k}w_{i}\mathcal{N}(\mu_{i},(\sigma_{i}+\bar{\sigma})^{2})$, and similarly for the backward policy, where $k$ is the number of Gaussians in the mixture.
+We schedule the value of $\bar{\sigma}$ so that it decreases during training according to an exponential-flat schedule:
+$\bar{\sigma}=\begin{cases}\bar{\sigma}_{0}\left(e^{-2je/(B/2)}-e^{-2e}\right)\;\;&j<B/2\\ 0\;\;&j\geq B/2,\end{cases}$ (13)
+where $j\in(1,\dots,B)$ is the batch number and $\bar{\sigma}_{0}=2$ is the initial noise, plotted in Figure 7.
+Note that, for the alanine dipeptide environment, the policy is a mixture of bivariate von Mises distributions and the noise $\bar{\sigma}$ is added to the concentration parameter $\kappa$, where concentration is related to standard deviation through $\sigma=\frac{1}{\kappa^{2}}$.
+Thompson sampling: We use $10$ heads with the bootstrapping probability parameter set to $p=0.3$.
+Local Search GFN: Local search samples trajectories using a forward-backward reconstruction method.
+First, forward policy samples actions to generate batch size $b$ on-policy trajectories.
+Then, the backward policy is used to re-sample the last $K$ steps of each trajectory.
+Finally, the forward policy is used to reconstruct the last $K$ steps, and trajectories in the batch with higher rewards from reconstruction replace the originals.
+We use $K=1$ for the line environment, alanine dipeptide environment and two-dimensional grid.
+We use $K=2$ and $K=3$ for the three and four dimensional (hyper)grids, respectively.
+Nested sampling: We use the Gradient Based Nested Sampling (GBNS) [34] with a publicly available implementation: https://github.com/Pablo-Lemos/GGNS.
+For each environment, GBNS generates samples from the reward distribution.
+During training, we alternate between two types of backpropagation: one fully on-policy, and the other using backward-sampled trajectories from the generated samples, guided by the current backward policy.
+Compute resources: Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU.
+D.2 Line environment details
+Parameterisation: The forward and backward policies are a mixture of three Gaussians.
+We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 256 hidden units per layer.
 We use the GELU activation function and dropout probability $0.2$ after each layer.
 This defines the torso of the MLP.
-Connecting from this common torso, the MLP has three single-layer, fully connected heads.
+Connecting from this common torso, the MLP has three single-layer, fully-connected heads.
 The first two heads have output dimension $9$ and parameterise the $3$ means ($\mu$), standard deviations ($\sigma$) and weights ($w$) of the mixture of Gaussians for the forward and backward policies respectively.
@@
 For the TB loss, we treat $\log Z_{\theta}$ as a separate learnable parameter.
-Replay buffer The replay buffer has capacity for $10^{4}$ trajectories.
-Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-3}$.
-When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
-Figure 8: Exponential noise schedule.
-Noisy exploration Noisy exploration is defined by adding an additional constant, $\bar{\sigma}$, to the standard deviations of the Gaussian distributions of the forward and backward policies.
-Specifically, the forward policy becomes $\hat{p}_{F}(s_{t},s_{t-1};\theta)=\sum_{i=1}^{3}w_{i}\mathcal{N}(\mu_{i},(\sigma_{i}+\bar{\sigma})^{2})$, and similarly for the backward policy.
-We schedule the value of $\bar{\sigma}$ so that it decreases during training according to an exponential-flat schedule:
-$\bar{\sigma}=\begin{cases}\bar{\sigma}_{0}\left(e^{-2je/(B/2)}-e^{-2e}\right)\;\;&j<B/2\\ 0\;\;&j\geq B/2,\end{cases}$ (13)
-where $j\in(1,\dots,B)$ is the batch number and $\bar{\sigma}_{0}=2$ is the initial noise, plotted in Figure 8.
-Thompson Sampling We use $10$ heads with the bootstrapping probability parameter set to $0.3$.
-MetaGFN We use $\Delta t=0.05$, $n=2$, $\beta=1$, $\gamma=2$, $w=0.15$, $\sigma=0.1$, $\epsilon=10^{-3}$.
-The domain of Adaptive Metadynamics is restricted to $[-5,23]$ and reflection conditions are imposed at the boundary.
+We use batch size $b=64$.
+It takes approximately $1$ hour to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
+MetaGFN: We use $\Delta t=0.05$, $n=2$, $\beta=1$, $\gamma=2$, $w=0.15$, $\sigma=0.1$, $\epsilon=10^{-3}$.
+The domain of Adapted Metadynamics is restricted to $[-5,23]$ and reflection conditions are imposed at the boundary.
 The bias and KDE potentials are stored on a uniform grid with grid spacing $0.01$.
 Initial metadynamics samples are drawn from a Gaussian distribution, mean 00 and variance $1$, and initial momenta from a Gaussian distribution, mean 00 and variance $0.5$.
-Training parameters In all experiments, we use batch size $b=64$ and for $B=10^{5}$ batches.
-We use a learning rate with a linear schedule, starting at $10^{-3}$ and finishing at 00.
-For the TB loss, we train the $\log Z_{\theta}$ with a higher initial learning rate of $10^{-1}$ (also linearly scheduled).
-For the STB loss, we use $\lambda=0.9$, a value that has worked well in the discrete setting Madan et al.
-[2022].
-We use the Adam optimiser with gradient clipping.
-For all loss functions, we clip the minimum log-reward signal at $-10$.
-This enables the model to learn despite regions of near-zero reward between the modes of $r(x)$.
-Evaluation The L1 error between the known reward distribution, $\rho(x)=r(x)/Z$, and the empirical on-policy distribution, denoted $\hat{\rho}(x)$, estimated by sampling $10^{4}$ on-policy trajectories and computing the empirical distribution over terminal states.
-Specifically, we compute
-$\text{error}=\frac{1}{2}\int_{-5}^{23}\left|\hat{\rho}(x)-\frac{r(x)}{Z}\right|\mathrm{d}x,$ (14)
-where the integral is estimated by a discrete sum with grid spacing $0.01$.
+Evaluation: The L1 error between the known reward distribution, $\rho(x)=r(x)/Z$, and the empirical on-policy distribution, denoted $\hat{\rho}(x)$, estimated by sampling $10^{4}$ on-policy trajectories and computing the empirical distribution over terminal states.
+Specifically, we compute $\text{error}=\frac{1}{2}\int_{-5}^{23}\left|\hat{\rho}(x)-\frac{r(x)}{Z}\right|\mathrm{d}x$, where the integral is estimated by a discrete sum with grid spacing $0.01$.
 Note that this error is normalised such that for all valid probability distributions $\hat{\rho}(x)$, we have $0\leq\text{error}\leq 1$.
-Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU.
-It takes approximately $1$ hour to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
-The additional computational expense of running Adaptive Metadynamics was negligible compared to the training time of the models.
-Figure 9: Forward policy and replay buffer distributions after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can consistently learn all three peaks.
-E.2 Results
-On-policy distributions Figure 9 shows the forward policy and replay buffer distributions (with bias sampling) after training for $10^{5}$ iterations with TB loss.
+On-policy distributions: Figure 9 shows the forward policy and replay buffer distributions (with bias sampling) after training for $10^{5}$ iterations with TB loss.
 MetaGFN is the only method that can uniformly populate the replay buffer and consistently learn all three peaks.
-Adaptive Metadynamics Figure 10(a) shows the L1 error between the density implied by the KDE potential and the true reward distribution during a typical training run.
+Adapted Metadynamics: Figure 10(a) shows the L1 error between the density implied by the KDE potential and the true reward distribution during a typical training run.
 Figure 10(b) shows the resulting $\hat{V}$ and $V_{\text{bias}}$ at the end of the training.
-By (1), Adaptive Metadynamics has fully explored the central peaks.
+By (1), Adapted Metadynamics has fully explored the central peaks.
 At (2), the third peak is discovered, prompting rapid adjustment of the KDE potential.
 By $2.5\times 10^{4}$ iterations, a steady state is reached and the algorithm is sampling the domain uniformly.
+Sensitivity analysis of hyperparameters: We compare MetaGFN performance for various freqRB and freqMD values, showing the mean and standard deviation of the final L1 error after $10^{5}$ batches (Figure 8).
+Intuitively, we expect that if freqRB is set too large, then MetaGFN focuses predominantly on the on-policy dynamics and may fail to learn to sample distant modes.
+By contrast, setting $\texttt{freqRB}=1$ means exclusively training on the replay buffer, without any on-policy training trajectories, which we would expect to lead to slow convergence.
+Therefore, an intermediate value of freqRB should be optimal.
+We found that setting $\texttt{freqRB}=2$ (giving an equal weighting to both on-policy and replay buffer trajectories) showed the lowest loss.
+By contrast, performance is not too sensitive to freqMD.
+freqMD should be set so that Adapted Metadynamics explores the reward landscape in a comparable time it takes to train the GFN to convergence, hence only the order of magnitude of freqMD is important.
+For concreteness, we choose $\texttt{freqMD}=10$ but the model performance is not too sensitive to this parameter.
+Figure 8: Mean and standard deviation L1 loss after $10^{5}$ training batches over $10$ repeats in the line environment.
+Blue line: changing freqRB while $\texttt{freqMD}=10$.
+Red line: changing freqMD while $\texttt{freqRB}=2$.
+Figure 9: Forward policy and replay buffer distributions after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can consistently learn all three peaks.
 Figure 10: (a) L1 error between $\hat{\rho}=\exp(-\beta\hat{V}(x))/Z$ and the reward distribution, $r(x)/Z$ (b) kde potential, bias potential, and positions of final samples after $2.5\times 10^{4}$ training iterations.
-Comparing MetaGFN Variants We consider three MetaGFN variants.
+Comparing MetaGFN variants: We consider three MetaGFN variants.
 The first variant, always backwards sample, regenerates the entire trajectory using the current backward policy when pulling from the replay buffer.
@@
 We observe that always backward sample is better than reuse initial backwards sample for all loss functions.
-For DB and TB losses, there is no evidence of any benefit of adding noise, whereas noise improves training for STB loss, performing very similarly to TB loss without noise.
+For DB and TB losses, there is no evidence for any benefit of adding noise, whereas noise improves training for STB loss, performing very similarly to TB loss without noise.
 Figure 11: Comparing MetaGFN variants.
-Appendix F Experiment details: grid environment
-Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 512 hidden units per layer, GELU activation function and dropout probability $0.2$, similar to the Line Environment.
-The first two heads have output dimension $24$ and parameterise the $8$ means, $12$ standard deviations and $4$ weights of the mixture of bivariate Gaussians.
-The third head has output dimension $1$ and parameterises the flow function $\hat{f}$.
-The means and standard deviations are mapped to the range $\mu\in(-15,15)$ and $\sigma\in(0.1,7)$ through sigmoid functions.
-Mixture weights are normalised with the softmax function.
-Replay buffer Identical to the Line Environment replay buffer, except trajectories are only stored when the terminal state’s reward exceeds $10^{-4}$.
-Noisy exploration We use the same noise profile as the Line Environment, equation (13). The noise $\bar{\sigma}$ is added equally to both dimensions of the standard deviation.
-Thompson sampling We use $10$ heads with the bootstrap probability parameter set to $0.3$.
-MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.35$, $n=3$, $\beta=1$, $\gamma=2$, $w=0.10$, $\sigma=2$.
-The bias and KDE potentials are stored on a uniform grid with grid spacing $0.075$.
-Initial samples are drawn from a Gaussian distribution, mean centred $(0,0)$, variance $\sigma^{2}=(1,1)$.
-Momenta are initialised to zero.
-Training parameters The same as for the Line Environment, see Appendix E.1.
-Evaluation The L1 error between the histogram of the terminal states of $10^{4}$ on-policy samples and the exact distribution is computed.
-$\text{error}=\frac{1}{2}\int_{-15}^{15}\int_{-15}^{15}\left|\hat{\rho}(x,y)-\frac{r(x,y)}{Z}\right|\mathrm{d}x\mathrm{d}y$, estimated by a discrete sum with grid spacing $0.075$.
-Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU.
-It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
-The additional computational expense of running Adaptive Metadynamics is less than $10\%$.
-Appendix G Experiment details: alanine dipeptide environment
-Computing the Free Energy Surface To obtain a ground-truth free energy surface (FES) in $\phi$-$\psi$ space, we ran a 250ns NPT well-tempered metadynamics MD simulation of alanine dipeptide at temperature 300K ($\beta=0.4009$), pressure 1bar with the TIP3P explicit water model Jorgensen et al.
-[1983].
-We used the PLUMMED plugin Bonomi et al.
-[2019] for OpenMM Eastman et al.
-[2017] with the AMBER14 force field Salomon-Ferrer et al.
-[2013].
-Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and $\hat{f}$ through three heads of an MLP with 3 hidden layers with 512 hidden units per layer, with GeLU activations and dropout probability $0.2$, similar to the Line Environment.
+D.3 Alanine dipeptide environment details
+Figure 12: The potential KDE $\hat{V}$ learnt using MetaGFN on the alanine dipeptide environment, concentration $\kappa=10$.
+Computing the free energy surface: To obtain a ground-truth free energy surface (FES) in $\phi$-$\psi$ space, we ran a 250ns NPT well-tempered metadynamics MD simulation of alanine dipeptide at temperature 300K ($\beta=0.4009$), pressure 1bar with the TIP3P explicit water model [47].
+We used the PLUMMED plugin [48] for OpenMM [49] with the AMBER14 force field [50].
+Parameterisation: The forward and backward policies are a mixture of three bivariate von Mises distributions.
+We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and $\hat{f}$ through three heads of an MLP with 3 hidden layers with 512 hidden units per layer, with GeLU activations and dropout probability $0.2$.
 The first two heads have output dimensions of $15$, parameterising the $6$ means, $6$ concentrations, and $3$ weights of the mixture of von Mises policy.
@@
 Mixture weights are normalised with the softmax function.
-Replay buffer The replay buffer has capacity for $10^{4}$ trajectories.
-Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-10}$.
-When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
-Noisy exploration We use the same noise profile as the Line Environment, equation (13).
-The noise $\bar{\sigma}$ is now added to the concentration parameter $\kappa$.
-Concentration is related to standard deviation through $\sigma=\frac{1}{\kappa^{2}}$.
-Thompson Sampling We use 10 heads with the bootstrap probability parameter set to 0.3.
-MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.01$, $n=2$, $\beta=0.4009$, $\gamma=0.1$, $w=10^{-5}$, $\kappa=10$, $\epsilon=10^{-6}$.
+We use batch size $b=64$.
+It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
+MetaGFN: We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.01$, $n=2$, $\beta=0.4009$, $\gamma=0.1$, $w=10^{-5}$, $\kappa=10$, $\epsilon=10^{-6}$.
 The bias and KDE potentials are stored on a uniform grid with grid spacing $0.1$.
 Initial samples are drawn from a Gaussian distribution, mean centred $P_{||}$, variance $\sigma^{2}=(0.1,0.1)$, and initial momenta from a Gaussian, mean $\mu=(0,0)$, variance $\sigma^{2}=(0.05,0.05)$.
-Training parameters The same as for the Line Environment, see Appendix E.1.
-Evaluation The L1 error of a histogram of on-policy samples, $\hat{\rho}(\phi,\psi)$, is computed via a two-dimensional generalisation of (14); $\text{error}=\frac{1}{2}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left|\hat{\rho}(\phi,\psi)-\frac{r(\phi,\psi)}{Z}\right|\mathrm{d}\phi\mathrm{d}\psi$, estimated by a discrete sum with grid spacing $0.1$.
-Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU.
+In Figure 12 we show the resulting learnt KDE potential with these parameters.
+Evaluation: The L1 error of a histogram of on-policy samples, $\hat{\rho}(\phi,\psi)$, is computed via a two-dimensional generalisation of (D.2); $\text{error}=\frac{1}{2}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left|\hat{\rho}(\phi,\psi)-\frac{r(\phi,\psi)}{Z}\right|\mathrm{d}\phi\mathrm{d}\psi$, estimated by a discrete sum with grid spacing $0.1$.
+D.4 Grid environments details
+Parameterisation: For the two-dimensional grid, we use a mixture of $4$ Gaussians.
+For the three and four dimensional (hyper)grids, we use a mixture of $2$ Gaussians (but use longer trajectories, see Section 4).
+We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 512 hidden units per layer, GELU activation function and dropout probability $0.2$.
+The means and standard deviations are mapped to the range $\mu\in(-15,15)$ and $\sigma\in(0.1,7)$ through sigmoid functions.
+Mixture weights are normalised with the softmax function.
+We use batch size $b=128$.
 It takes approximately 10 hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
-The additional computational expense of running Adaptive Metadynamics was less than $10\%$.
+MetaGFN: We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.35$, $n=3$, $\beta=1$, $\gamma=2$, $w=0.10$, $\sigma=2$.
+The bias and KDE potentials are stored on a uniform grid with grid spacing $0.075/0.3/0.6$ (2, 3, and 4 dimensions respectively).
+Initial samples are drawn from a Gaussian distribution, mean centred at the origin, isotropic variance $\sigma^{2}=1$.
+Momenta are initialised to zero.
+Evaluation: The L1 error between the histogram of the terminal states of $10^{4}$ on-policy samples and the exact distribution is computed, e.g.
+for the 2D grid, $\text{error}=\frac{1}{2}\int_{-15}^{15}\int_{-15}^{15}\left|\hat{\rho}(x,y)-\frac{r(x,y)}{Z}\right|\mathrm{d}x\mathrm{d}y$, estimated by a discrete sum with grid spacing $0.075$.